Linux NET-3-HOWTO, Linux Networking.

Terry Dawson, VK2KTJ, terry@perf.no.itg.telstra.com.au v1.2, 20 August 1997 The Linux Operating System boasts kernel based networking support written almost entirely from scratch. The performance of the tcp/ip implementation in recent kernels makes it a worthy alternative to even the best of its peers. This document aims to describe how to install and configure the Linux networking software and associated tools.

Contents

1 Changes from the previous version 4

2 Introduction. 5 2.1 Feedback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 How to use this HOWTO document (NET-3-HOWTO-HOWTO ?). 5

4 General Information about Linux Networking. 6 4.1 A brief history of Linux Networking Kernel Development. . . . . . . . . . . . . . . . . . . . . 6 4.2 Where to get other information about Linux Networking. . . . . . . . . . . . . . . . . . . . . 8 4.3 Where to get some non-linux-specific network information. . . . . . . . . . . . . . . . . . . . . 9

5 Generic Network Configuration Information. 10

5.1 What do I need to start ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.1.1 Current Kernel source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.1.2 Current Network tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.1.3 Network Application Programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.1.4 Addresses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5.2 Where should I put the configuration commands ? . . . . . . . . . . . . . . . . . . . . . . . . 14 5.3 Creating your network interfaces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.4 Configuring a network interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.5 Configuring your Name Resolver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5.5.1 What’s in a name ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.5.2 What information you will need. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.5.3 /etc/resolv.conf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.5.4 /etc/host.conf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.5.5 /etc/hosts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

5.6 Configuring your loopback interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.7 Routing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

5.7.1 So what does the routed program do ? . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.8 Configuring your network servers and services. . . . . . . . . . . . . . . . . . . . . . . . . . . 23

5.8.1 /etc/services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 5.8.2 /etc/inetd.conf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

5.9 Other miscellaneous network related configuration files. . . . . . . . . . . . . . . . . . . . . . 32 5.9.1 /etc/protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 5.9.2 /etc/networks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

5.10 Network Security and access control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 5.10.1 /etc/ftpusers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 5.10.2 /etc/securetty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 5.10.3 The tcpd hosts access control mechanism. . . . . . . . . . . . . . . . . . . . . . . . . . 34 5.10.4 /etc/hosts.equiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 5.10.5 Configure your ftp daemon properly. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 5.10.6 Network Firewalling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 5.10.7 Other suggestions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

6 Network Technology Specific Information. 36 6.1 ARCNet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 6.2 Appletalk (AF APPLETALK) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

6.2.1 Configuring the Appletalk software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.2.2 Exporting a Linux filesystems via Appletalk. . . . . . . . . . . . . . . . . . . . . . . . 38 6.2.3 Sharing your Linux printer across Appletalk. . . . . . . . . . . . . . . . . . . . . . . . 39 6.2.4 Starting the appletalk software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.2.5 Testing the appletalk software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.2.6 Caveats of the appletalk software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.2.7 More information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

6.3 ATM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.4 AX25 (AF AX25) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.5 DECNet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.6 EQL - multiple line traffic equaliser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.7 Ethernet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.8 FDDI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 6.9 Frame Relay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 6.10 IP Accounting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 6.11 IP Aliasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 6.12 IP Firewall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 6.13 IPIP Encapsulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.13.1 A tunneled network configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 6.13.2 A tunneled host configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

6.14 IPX (AF IPX) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 6.15 IPv6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.16 ISDN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.17 IP Masquerade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 6.18 IP Transparent Proxy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.19 Mobile IP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 6.20 Multicast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 6.21 NAT - Network Address Translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 6.22 NetRom (AF NETROM) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 6.23 PLIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 6.24 PPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

6.24.1 Maintaining a permanent connection to the net with pppd. . . . . . . . . . . . . . . . 62 6.25 Rose protocol (AF ROSE) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 6.26 SAMBA - ‘NetBEUI’, ‘NetBios’ support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.27 SLIP client . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

6.27.1 dip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.27.2 slattach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.27.3 When do I use which ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.27.4 Static SLIP server with a dialup line and DIP. . . . . . . . . . . . . . . . . . . . . . . 65 6.27.5 Dynamic SLIP server with a dialup line and DIP. . . . . . . . . . . . . . . . . . . . . . 65 6.27.6 Using DIP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.27.7 Permanent SLIP connection using a leased line and slattach. . . . . . . . . . . . . . . 68 6.28 SLIP server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

6.28.1 Slip Server using sliplogin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

6.28.2 Slip Server using dip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

6.28.3 SLIP server using the dSLIP package. . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

6.29 STRIP support (Starmode Radio IP) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

6.30 Token Ring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

6.31 X.25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

6.32 WaveLan Card . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

7 Cables and Cabling 78

7.1 Serial NULL Modem cable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

7.2 Parallel port cable (PLIP cable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

7.3 10base2 (thin coax) Ethernet Cabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

7.4 Twisted Pair Ethernet Cable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

8 Glossary of Terms used in this document. 80

9 Linux for an ISP ? 81

10 Acknowledgements 82

11 Copyright. 82

1 Changes from the previous version

Additions:
        Reference to PLIP-mini-HOWTO - thanks Claes
        IP NAT - Network Address Translation
Corrections/Updates:
        Many corrections from Alessandro Rubini - thanks!
        Updated Larry Stefani's email address - thanks Larry
        Corrected ftp.linux.uk.org nettools location - thanks Ron
        Corrected incorrect route command - thanks John
        More broken route commands! - thanks Jean-Pierre
        IPv6 addresses are 16 bytes not 32, oops - thanks Erez
ToDo:

        Add traffic shaper
        Describe new routing algorithm
        Add IPv6 kernel compile options
        Describe /proc/sys/net/* entries.
        WanRouter device

2 Introduction.

The original NET-FAQ was written by Matt Welsh and I to answer frequently asked questions about networking for Linux at a time before the Linux Documentation Project had formally started. It covered the very early development versions of the Linux Networking Kernel. The NET-2-HOWTO superceded the NET-FAQ and was one of the original LDP HOWTO documents, it covered what was called version 2 and later version 3 of the Linux kernel Networking software. This document in turn supercedes it and relates only to version 3 of the Linux Networking Kernel.

Previous versions of this document became quite large because of the enormous amount of material that fell within its scope. To help reduce this problem a number of HOWTO’s dealing with specific networking topics have been produced. This document will provide pointers to them where relevant and cover those areas not yet covered by other documents.

2.1 Feedback

I always appreciate feedback and especially value contributions. Please direct any feedback or contributions

to me by email <mailto:terry@perf.no.itg.telstra.com.au>.

3 How to use this HOWTO document (NET-3-HOWTO-HOWTO ?).

The format of this document is differs from earlier versions. I’ve now regrouped the sections so that there is informative material at the beginning which you can skip if you are not interested, generic material next which you must ensure you understand before proceeding to the technology specific sections in the rest of the document.

Read the generic sections These sections apply to every, or nearly every, technology described later and so are very important for you to understand.

Consider your network You should know how your network is, or will be, designed and exactly what hardware and technology types you will be implementing.

Read the technology specific sections related to your requirements When you know what you want you can address each component in turn. These sections cover only details specific to a particular technology.

Do the configuration work You should actually try to configure your network and take careful note of any problems you have.

Look for further help if needed If you experience problems that this document does not help you to resolve then read the section related to where to get help or where to report bugs.

Have fun!

Networking is fun, enjoy it.

4 General Information about Linux Networking.

4.1 A brief history of Linux Networking Kernel Development.

Developing a brand new kernel implementation of the tcp/ip protocol stack that would perform as well as existing implementations was not an easy task. The decision not to port one of the existing implementations was made at a time when there was some uncertainty as to whether the existing implementations may become encumbered by restrictive copyrights because of the court case put by U.S.L. and when there was a lot of fresh enthusiasm for doing it differently and perhaps even better than had already been done.

The original volunteer to lead development of the kernel network code was Ross Biro <biro@yggdrasil.com>. Ross produced a simple and incomplete but mostly usable implementation set of routines that were complemented by an ethernet driver for the WD-8003 network interface card. This was enough to get many people testing and experimenting with the software and some people even managed to connect machines in this configuration to live internet connections. The pressure within the Linux community driving development for networking support was building and eventually the cost of a combination of some unfair pressure applied to Ross and his own personal commitments outweighed the benefit he was deriving and he stepped down as lead developer. Ross’s efforts in getting the project started and accepting the responsibility for actually producing something useful in such controversial circumstances were what catalysed all future work and were therefore an essential component of the success of the current product.

Orest Zborowski <obz@Kodak.COM> produced the original BSD socket programming interface for the Linux kernel. This was a big step forward as it allowed many of the existing network applications to be ported to linux without serious modification.

Somewhere about this time Laurence Culhane <loz@holmes.demon.co.uk> developed the first drivers for Linux to support the SLIP protocol. These enabled many people who did not have access to Ethernet networking to experiment with the new networking software. Again, some people took this driver and pressed it into service to connect them to the Internet. This gave many more people a taste of the possibilities that could be realised if Linux had full networking support and grew the number of users actively using and experimenting with the networking software that existed.

One of the people that had also been actively working on the task of building networking support was Fred van Kempen <waltje@uwalt.nl.mugnet.org>. After a period of some uncertainty following Ross’s resignation from the lead developer position Fred offered his time and effort and accepted the role essentially unopposed. Fred had some ambitious plans for the direction that he wanted to take the Linux networking software and he set about progressing in those directions. Fred produced a series of networking code called the ‘NET-2’ kernel code (the ‘NET’ code being Ross’s) which many people were able to use pretty much usefully. Fred formally put a number of innovations on the development agenda, such as the dynamic device interface, Amateur Radio AX.25 protocol support and a more modularly designed networking implementation. Fred’s NET-2 code was used by a fairly large number of enthusiasts, the number increasing all the time as word spread that the software was working. The networking software at this time was still a large number of patches to the standard release of kernel code and was not included in the normal release. The NET-FAQ and subsequent NET-2-HOWTO’s described the then fairly complex procedure to get it all working. Fred’s focus was on developing innovations to the standard network implementations and this was taking time. The community of users was growing impatient for something that worked reliably and satisfied the 80% of users and, as with Ross, the pressure on Fred as lead developer rose.

Alan Cox <iialan@www.uk.linux.org> proposed a solution to the problem designed to resolve the situation. He proposed that he would take Fred’s NET-2 code and debug it, making it reliable and stable so that it would satisfy the impatient user base while relieving that pressure from Fred allowing him to continue his work. Alan set about doing this, with some good success and his first version of Linux networking code was called ‘Net-2D(ebugged)’. The code worked reliably in many typical configurations and the user base was happy. Alan clearly had ideas and skills of his own to contribute to the project and many discussions relating to the direction the NET-2 code was heading ensued. There developed two distinct schools within the Linux networking community, one that had the philosophy of ‘make it work first, then make it better’ and the other of ‘make it better first’. Linus ultimately arbitrated and offered his support to Alan’s development efforts and included Alan’s code in the standard kernel source distribution. This placed Fred in a difficult position. Any continued development would lack the large user base actively using and testing the code and this would mean progress would be slow and difficult. Fred continued to work for a short time and eventually stood down and Alan came to be the new leader of the Linux networking kernel development effort.

Donald Becker <becker@cesdis.gsfc.nasa.gov> soon revealed his talents in the low level aspects of networking and produced a huge range of ethernet drivers, nearly all of those included in the current kernels were developed by Donald. There have been other people that have made significant contributions, but Donald’s work is prolific and so warrants special mention.

Alan continued refining the NET-2-Debugged code for some time while working on progressing some of the matters that remained unaddressed on the ‘TODO’ list. By the time the Linux 1.3.* kernel source had grown its teeth the kernel networking code had migrated to the NET-3 release on which current versions are based. Alan worked on many different aspects of the networking code and with the assistance of a range of other talented people from the Linux networking community grew the code in all sorts of directions. Alan produced dynamic network devices and the first standard AX.25 and IPX implementations. Alan has continued tinkering with the code, slowly restructuring and enhancing it to the state it is in today.

PPP support was added by Michael Callahan <callahan@maths.ox.ac.uk> and Al Longyear <longyear@netcom.com> this too was critical to increasing the number of people actively using linux for networking.

Jonathon Naylor <jsn@cs.nott.ac.uk> has contributed by significantly enhancing Alan’s AX.25 code, adding NetRom and Rose protocol support. The AX.25/NetRom/Rose support itself is quite significant, because no other operating system can boast standard native support for these protocols beside Linux.

There have of course been hundreds of other people who have made significant contribution to the development of the Linux networking software. Some of these you will encounter later in the technology specific sections, other people have contributed modules, drivers, bug-fixes, suggestions, test reports and moral support. In all cases each can claim to have played a part and offered what they could. The Linux kernel networking code is an excellent example of the results that can be obtained from the Linux style of anarchic development, if it hasn’t yet surprised you, it is bound to soon enough, the development hasn’t stopped.

4.2 Where to get other information about Linux Networking.

There are a number of places where you can find good information about Linux networking.

Alan Cox, the current maintainer of the Linux kernel networking code maintains a world wide web page that contains highlights of current and new developments in linux Networking at: www.uk.linux.org

<http://www.uk.linux.org/NetNews.html>.

Another good place is a book written by Olaf Kirch entitled the Network Administrators Guide. It is a work of the Linux Documentatation Project <http://sunsite.unc.edu/LDP/> and you can read it interactively at Network Administrators Guide HTML version

<http://sunsite.unc.edu/LDP/LDP/nag/nag.html> or you can obtain it in various formats by ftp from
the sunsite.unc.edu LDP ftp archive <ftp://sunsite.unc.edu/pub/Linux/docs/LDP/network-guide/>.

Olaf’s book is quite comprehensive and provides a good high level overview of network configuration under linux.

There is a newsgroup in the Linux news heirarchy dedicated to networking and related matters, it is:

comp.os.linux.networking <news:comp.os.linux.networking>

There is a mailing list to which you can subscribe where you may ask questions relating to Linux networking. To subscribe you should send a mail message:

     To: majordomo@vger.rutgers.edu
     Subject: anything at all
     Message:
     subscribe linux-net

On the various IRC networks there are often #linux channels on which people will be able to answer questions on linux networking.

Please remember when reporting any problem to include as much relevant detail about the problem as you can. Specifically you should specify the versions of software that you are using, especially the kernel version, the version of tools such as pppd or dip and the exact nature of the problem you are experiencing. This means taking note of the exact syntax of any error messages you receive and of any commands that you are issuing.

4.3 Where to get some non-linux-specific network information.

If you are after some basic tutorial information on tcp/ip networking generally, then I recommend you take a look at the following documents:

tcp/ip introduction

     this document comes as both a text version <ftp://athos.rutgers.edu/runet/tcp-ip-intro.doc>
     and a postscript version <ftp://athos.rutgers.edu/runet/tcp-ip-intro.ps>.

tcp/ip administration

     this document comes as both a text version <ftp://athos.rutgers.edu/runet/tcp-ip-admin.doc>
     and a postscript version <ftp://athos.rutgers.edu/runet/tcp-ip-admin.ps>.

If you are after some more detailed information on tcp/ip networking then I highly recommend:

     "Internetworking with TCP/IP"
     by Douglas E. Comer
     ISBN 0-13-474321-0
     Prentice Hall publications.

If you are wanting to learn about how to write network applications in a Unix compatible environment then I also highly recommend:

     "Unix Network Programming"
     by W. Richard Stevens
     ISBN 0-13-949876-1
     Prentice Hall publications.

You might also try the comp.protocols.tcp-ip <news:comp.protocols.tcp-ip> newsgroup.

An important source of specific technical information relating to the Internet and the tcp/ip suite of protocols are RFC’s. RFC is an acronym for ‘Request For Comment’ and is the standard means of submitting and documenting Internet protocol standards. There are many RFC repositories. Many of these sites are ftp sites and other provide World Wide Web access with an associated search engine that allows you to search the RFC database for particular keywords.

One possible source for RFC’s is at: Nexor RFC database

<http://pubweb.nexor.co.uk/public/rfc/index/rfc.html>.

5 Generic Network Configuration Information.

The following subsections you will pretty much need to know and understand before you actually try to configure your network. They are fundamental principles that apply regardless of the exact nature of the network you wish to deploy.

5.1 What do I need to start ?

Before you start building or configuring your network you will need some things. The most important of these are:

5.1.1 Current Kernel source.

Because the kernel you are running now might not yet have support for the network types or cards that you wish to use you will probably need the kernel source so that you can recompile the kernel with the appropriate options.

You can always obtain the latest kernel source from: ftp.funet.fi

<ftp://ftp.funet.fi/pub/Linux/PEOPLE/Linus/v2.0>.

Normally the kernel source will be untarred into the /usr/src/linux directory. For information on how to apply patches and build the kernel you should read the Kernel-HOWTO <Kernel-HOWTO.html>. For information on how to configure kernel modules you should read the Module-HOWTO

<Module-HOWTO.html>.

Unless specifically stated otherwise, I recommend you stick with the standard kernel release (the one with the even number as the second digit in the version number). Development release kernels (the ones with the odd second digit) may have structural or other changes that may cause problems working with the other software on your system. If you are uncertain that you could resolve those sorts of problems in addition to the potential for there being other software errors, then don’t use them.

5.1.2 Current Network tools.

The network tools are the programs that you use to configure linux network devices. These tools allow you to assign addresses to devices and configure routes for example.

Most modern linux distributions are supplied with the network tools, so if you have installed from a distribution and haven’t yet installed the network tools then you should do so.

If you haven’t installed from a distribution then you will need to source and compile the tools yourself. This isn’t difficult.

The network tools are now maintained by Bernd Eckenfels and are available at: ftp.inka.de

<ftp://ftp.inka.de/pub/comp/Linux/networking/NetTools/> and are mirrored at: ftp.uk.linux.org
<ftp://ftp.uk.linux.org/pub/linux/Networking/base/>.

Be sure to choose the version that is most appopriate for the kernel you wish to use and follow the instructions in the package to install.

To install and configure the version current at the time of the writing you need do the following:

#
# cd /usr/src
# tar xvfz net-tools-1.33.tar.gz
# cd net-tools-1.33
# make config
# make
# make install
#

Additionally, if you intend configuring a firewall or using the IP masquerade feature you will require the ipfwadm command. The latest version of it may be obtained from: ftp.xos.nl <ftp:/ftp.xos.nl/pub/linux/ipfwadm>. Again there are a number of versions available. Be sure to pick the version that most closely matches your kernel.

To install and configure the version current at the time of the writing you need do the following:

#
# cd /usr/src
# tar xvfz ipfwadm-2.3.0.tar.gz
# cd ipfwadm-2.3.0
# make
# make install
#

5.1.3 Network Application Programs.

The network application programs are programs such as telnet and ftp and their respective server programs. David Holland <dholland@hcs.harvard.edu> now manages a distribution of the most common of these.

You may obtain it from: ftp.uk.linux.org <ftp://ftp.uk.linux.org/pub/linux/Networking/base>.

To install and configure the version current at the time of the writing you need do the following:

#
# cd /usr/src
# tar xvfz /pub/net/NetKit-B-0.08.tar.gz
# cd NetKit-B-0.08
# more README
# vi MCONFIG
# make
# make install
#

5.1.4 Addresses.

Internet Protocol Addresses are composed of four bytes. The convention is to write addresses in what is called ‘dotted decimal notation’. In this form each byte is converted to a decimal number (0-255) dropping any leading zero’s unless the number is zero and written with each byte seperated by a ‘.’ character. By convention each interface of a host or router has an IP address. It is legal for the same IP address to be used on each interface of a single machine in some circumstances but usually each interface will have its own address.

Internet Protocol Networks are contiguous sequences of IP addresses. All addresses within a network have a number of digits within the address in common. The portion of the address that is common amongst all addresses within the network is called the ‘network portion’ of the address. The remaining digits are called the ‘host portion’. The number of bits that are shared by all addresses within a network is called the netmask and it is role of the netmask to determine which addresses belong to the network it is applied to and which don’t. For example, consider the following:

-----------------  ---------------
Host Address       192.168.110.23
Network Mask       255.255.255.0
Network Portion    192.168.110.
Host portion                  .23
-----------------  ---------------
Network Address    192.168.110.0
Broadcast Address  192.168.110.255
-----------------  ---------------

Any address that is ’bitwise anded’ with its netmask will reveal the address of the network it belongs to. The network address is therefore always the lowest numbered address within the range of addresses on the network and always has the host portion of the address coded all zeroes.

The broadcast address is a special address that every host on the network listens to in addition to its own unique address. This address is the one thhat datagrams are sent to if every host on the network is meant to receive it. Certain types of data like routing information and warning messages are transmitted to the broadcast address so that every host on the network can receive it simultaneously. There are two commonly used standards for what the broadcast address should be. The most widely accepted one is to use the highest possible address on the network as the broadcast address. In the example above this would be 192.168.110.255. For some reason other sites have adopted the convention of using the network address as the broadcast address. In practice it doesn’t matter very much which you use but you must make sure that every host on the network is configured with the same broadcast address.

For administrative reasons some time early in the development of the IP protocol some arbitrary groups of addresses were formed into networks and these networks were grouped into what are called classes. These classes provide a number of standard size networks that could be allocated. The ranges allocated are:

----------------------------------------------------------
| Network | Netmask       | Network Addresses            |

| Class   |               |                              |
----------------------------------------------------------
|    A    | 255.0.0.0     | 0.0.0.0    - 127.255.255.255 |
|    B    | 255.255.0.0   | 128.0.0.0  - 191.255.255.255 |
|    C    | 255.255.255.0 | 192.0.0.0  - 223.255.255.255 |
|Multicast| 240.0.0.0     | 224.0.0.0  - 239.255.255.255 |
----------------------------------------------------------

What addresses you should use depends on exactly what it is that you are doing. You may have to use a combination of the following activities to get all the addresses you need:

Installing a linux machine on an existing IP network If you wish to install a linux machine onto an existing IP network then you should contact whoever administers the network and ask them for the following information:

• Host IP Address

• IP network address

• IP broadcast address

• IP netmask

• Router address

• Domain Name Server Address

You should then configure your linux network device with those details. You can not make them up and expect your configuration to work.

Building a brand new network that will never connect to the Internet If you are building a private network and you never intend that network to be connected to the Internet then you can choose whatever addresses you like. However, for safety and consistency reasons there have been some IP network addresses that have been reserved specifically for this purpose. These are specified in RFC1597 and are as follows:

-----------------------------------------------------------
|         RESERVED PRIVATE NETWORK ALLOCATIONS            |
-----------------------------------------------------------
| Network | Netmask       | Network Addresses             |
| Class   |               |                               |
-----------------------------------------------------------
|    A    | 255.0.0.0     | 10.0.0.0    - 10.255.255.255  |
|    B    | 255.255.0.0   | 172.16.0.0  - 172.31.255.255  |
|    C    | 255.255.255.0 | 192.168.0.0 - 192.168.255.255 |
-----------------------------------------------------------

You should first decide how large you want your network to be and then choose as many of the addresses as you require.

5.2 Where should I put the configuration commands ?

There are a few different approaches to Linux system boot procedures. After the kernel boots, it always executes a program called ‘init’. The init program then reads its configuration file called /etc/inittab and commences the boot process. There are a few different flavours of init and it is this variation that is the largest cause of variation between distributions or machines.

Usually the /etc/inittab file contains an entry looking something like:

     si::sysinit:/etc/init.d/boot

This line specifies the name of the shell script file that actually manages the boot sequence. This file is somewhat equivalent to the AUTOEXEC.BAT file in MS-DOS.

There are usually other scripts that are called by the boot script and often the network is configured within one of many of these.

The following table may be used as a guide for your system:

-------------------------------------------------------------------------------
Distrib. |Interface Config/Routing                    |Server Initialisation
-------------------------------------------------------------------------------
Debian   |/etc/init.d/network                         |/etc/init.d/netbase
         |                                            |/etc/init.d/netstd_init
         |                                            |/etc/init.d/netstd_nfs
         |                                            |/etc/init.d/netstd_misc
-------------------------------------------------------------------------------
Slackware|/etc/rc.d/rc.inet1                          |/etc/rc.d/rc.inet2
-------------------------------------------------------------------------------
RedHat   |/etc/sysconfig/network-scripts/ifup-{$<$}ifname>|/etc/rc.d/init.d/network
-------------------------------------------------------------------------------

Most modern distributions include a program that will allow you to configure many of the common sorts of network interfaces. If you have one of these then you should see if it will do what you want before attempting a manual configuration.

     -----------------------------------------
     Distrib   | Network configuration program
     -----------------------------------------
     RedHat    | /sbin/netcfg
     Slackware | /sbin/netconfig
     -----------------------------------------

5.3 Creating your network interfaces.

In many Unix operating systems the network devices have appearances in the /dev directory. This is not so in Linux. In Linux the network devices are created dynamically in software and thus do not require device files to be present.

In the majority of cases the network devices is automatically created by the device driver while it is initialising and has located your hardware. For example, the ethernet device driver creates eth[0..n] interfaces sequentially as it locates your ethernet hardware. The first ethernet card found becomes eth0, the second eth1 etc.

In some cases though, notably slip and ppp, the network devices are created through the action of some user program. The same sequential device numbering applies, but the devices are not created automatically at boot time. The reason for this is that unlike ethernet devices, the number of active slip or ppp devices may vary during the uptime of the machine. These cases will be covered in more detail in later sections.

5.4 Configuring a network interface.

When you have all of the programs you need and your address and network information you can configure your network interfaces. When we talk about configuring a network interface we are talking about the process of assigning appropriate addresses to a network device and to setting appropriate values for other configurable paramaters of a network device. The program most commonly used to do this is the ifconfig (interface configure) command.

Typically you would use a command similar to the following:

     # ifconfig eth0 192.168.0.1 netmask 255.255.255.0 up

In this case I’m configuring an ethernet interface ‘eth0’ with the IP address ‘192.168.0.1’ and a network mask of ‘255.255.255.0’. The ‘up’ that trails the command tells the interface that it should become active.

The kernel assumes certain defaults when configuring interfaces. For example, you may specify the network address and broadcast address for an interface, but if you don’t, as in my example above, then the kernel will make reasonable guesses as to what they should be based on the netmask you supply and if you don’t supply a netmask then on the network class of the IP address configured. In my example the kernel would assume that it is a class-C network being configured on the interface and configure a network address of ‘192.168.0.0’ and a broadcast address of ‘192.168.0.255’ for the interface.

There are many other options to the ifconfig command. The most important of these are:

up this option activates an interface.

down this option deactivates an interface.

[- arp] this option enables or disables use of the address resolution protocol on this interface

[- allmulti] this option enables or disables the reception of all hardware multicast packets. Hardware multicast enables groups of hosts to receive packets addressed to special destinations. This may be of importance if you are using applications like desktop videoconferencing but is normally not used. mtu N

this parameter allows you to set the MTU of this device.

netmask addr this parameter allows you to set the network mask of the network this device belongs to.

irq addr this parameter only works on certain types of hardware and allows you to set the IRQ of the hardware of this device.

[- broadcast [addr]] this parameter allows you to enable and set the accepting of datagrams destined to the broadcast address, or to disable reception of these datagrams.

[- pointopoint [addr]] this parameter allows you to set the address of the machine at the remote end of a point to point link such as for slip or ppp.

hw <type¿ <addr¿ this parameter allows you to set the hardware address of certain types of network devices. This is not often useful for ethernet, but is useful for other network types such as AX.25.

You may use the ifconfig command on any network interface. Some user programs such as pppd and dip automatically configure the network devices as they create them, so manual use of ifconfig is unnecessary.

5.5 Configuring your Name Resolver.

The ‘Name Resolver’ is a part of the linux standard library. Its prime function is to provide a service to convert human-friendly hostnames like ‘ftp.funet.fi’ into machine friendly IP addresses such as

128.214.248.6.

5.5.1 What’s in a name ?

You will probably be familiar with the appearance of Internet host names, but may not understand how they are constructed, or deconstructed. Internet domain names are heirarchial in nature, that is, they have a tree-like structure. A ‘domain’ is a family, or group of names. A ‘domain’ may be broken down into ‘subdomain’. A ‘toplevel domain’ is a domain that is not a subdomain. The Top Level Domains are specified in RFC-920. Some examples of the most common top level domains are:

COM Commercial Organisations

EDU Educational Organisations GOV

Government Organisations

MIL Millitary Organisations

ORG Other organisations

Country Designator these are two letters codes that represent a particular country.

Each of these top level domains has subdomains. The top level domains based on country name are used next broken down into subdomains based on the com, edu, gov, mil and org domains. So for example you end up with: com.au and gov.au for commercial and government organisations in Australia. For historical reasons most domains belonging to one of the non-country based top level domains are for organisations within the United States, although the United States also has its own country code ‘.us’.

The next level of division usually represents the name of the organisation. Further subdomains vary in nature, often the next level of subdomain is based on the departmental structure of the organisation but it may be based on any criterion considered reasonable and meaningful by the network adminstrators for the organisation.

The very left-most portion of the name is always the unique name assigned to the host machine and is called the ‘hostname’, the portion of the name to the right of the hostname is called the ‘domainname’ and the complete name is called the ‘Fully Qualified Domain Name’.

To use my own email host as an example, the fully qualified domain name is ‘perf.no.itg.telstra.com.au’. This means that the host name is ‘perf’ and the domain name is ‘no.itg.telstra.com.au’. The domain name is based on a top level domain based on my country, Australia and as my email address belongs to a commercial organisation we have ‘.com’ as the next level domain. The name of the company is (was) ‘telstra’ and our internal naming structure is based on organisational structure, in my case, my machine belongs to the Information Technology Group, Network Operations section.

5.5.2 What information you will need.

You will need to know what domain your hosts name will belong to. The name resolver software provides this name translation service by making requests to a ‘Domain Name Server’, so you will need to know the IP address of a local nameserver that you can use.

There are three files you need to edit, I’ll cover each of these in turn.

5.5.3 /etc/resolv.conf

The /etc/resolv.conf is the main configuration file for the name resolver code. Its format is quite simple. It is a text file with one keyword per line. There are three keywords typically used, they are: domain

this keyword specifies the local domain name.

search this keyword specifies a list of alternate domain names to search for a hostname

nameserver this keyword, which may be used many times, specifies an IP address of a domain name server to query when resolving names

An example /etc/resolv.conf might look something like:

     domain maths.wu.edu.au
     search maths.wu.edu.au wu.edu.au
     nameserver 192.168.10.1
     nameserver 192.168.12.1

This example specifies that the default domain name to append to unqualified names (ie hostnames supplied without a domain) is maths.wu.edu.au and that if the host is not found in that domain to also try the wu.edu.au domain directly. Two nameservers entry are supplied, each of which may be called upon by the name resolver code to resolve the name.

5.5.4 /etc/host.conf

The /etc/host.conf file is where you configure some items that govern the behaviour of the name resolver code. The format of this file is described in detail in the ‘resolv+’ man page. In nearly all circumstances the following example will work for you:

     order hosts,bind
     multi on

This configuration tells the name resolver to check the /etc/hosts file before attempting to query a nameserver and to return all valid addresses for a host found in the /etc/hosts file instead of just the first.

5.5.5 /etc/hosts

The /etc/hosts file is where you put the name and IP address of local hosts. If you place a host in this file then you do not need to query the domain name server to get its IP Address. The disadvantage of doing this is that you must keep this file up to date yourself if the IP address for that host changes. In a well managed system the only hostnames that usually appear in this file are an entry for the loopback interface and the local hosts name.

     # /etc/hosts
     127.0.0.1      localhost loopback
     192.168.0.1    this.host.name

You may specify more than one host name per line as demonstrated by the first entry, which is a standard entry for the loopback interface.

5.6 Configuring your loopback interface.

The ‘loopback’ interface is a special type of interface that allows you to make connections to yourself. There are various reasons why you might want to do this, for example, you may wish to test some network software without interfering with anybody else on your network. By convention the IP address ‘127.0.0.1’ has been assigned specifically for loopback. So no matter what machine you go to, if you open a telnet connection to 127.0.0.1 you will always reach the local host.

Configuring the loopback interface is simple and you should ensure you do.

     # ifconfig lo 127.0.0.1
     # route add -host 127.0.0.1 lo

We’ll talk more about the route command in the next section.

5.7 Routing.

Routing is a big topic. It is easily possible to write large volumes of text about it. Most of you will have fairly simple routing requirements, some of you will not. I will cover some basic fundamentals of routing only. If you are interested in more detailed information then I suggest you refer to the references provided at the start of the document.

Let’s start with a definition. What is IP routing ? Here is one that I’m using:

IP Routing is the process by which a host with multiple network connections decides where to deliver IP datagrams it has received.

It might be useful to illustrate this with an example. Imagine a typical office router, it might have a PPP link off the Internet, a number of ethernet segments feeding the workstations and another PPP link off to another office. When the router receives a datagram on any of its network connections, routing is the mechanism that it uses to determine which interface it should send the datagram to next. Simple hosts also need to route, all Internet hosts have two network devices, one is the loopback interface described above and the other is the one it uses to talk to the rest of the network, perhaps an ethernet, perhaps a PPP or SLIP serial interface.

Ok, so how does routing work ? Each host keeps a special list of routing rules, called a routing table. This table contains rows which typically contain at least three fields, the first is a destination address, the second is the name of the interface to which the datagram is to be routed and the third is optionally the IP address of another machine which will carry the datagram on its next step through the network. In linux you can see this table by using the following command:

     # cat /proc/net/route

or by using either of the following commands:

     # /sbin/route -n
     # /bin/netstat -r

The routing process is fairly simple: an incoming datagram is received, the destination address (who it is for) is examined and compared with each entry in the table. The entry that best matches that address is selected and the datagram is forwarded to the specified interface. If the gateway field is filled then the datagram is forwarded to that host via the specified interface, otherwise the destination address is assumed to be on the network supported by the interface.

To manipulate this table a special command is used. This command takes command line arguments and converts them into kernel system calls that request the kernel to add, delete or modify entries in the routing table. The command is called ‘route’.

A simple example. Imagine you have an ethernet network. You’ve been told it is a class-C network with an address of 192.168.1.0. You’ve been supplied with an IP address of 192.168.1.10 for your use and have been told that 192.168.1.1 is a router connected to the Internet.

The first step is to configure the interface as described earlier. You would use a command like:

     # ifconfig eth0 192.168.1.10 netmask 255.255.255.0 up

You now need to add an entry into the routing table to tell the kernel that datagrams for all hosts with addresses that match 192.168.1.* should be sent to the ethernet device. You would use a command similar to:

     # route add -net 192.168.1.0 netmask 255.255.255.0 eth0

Note the use of the ‘-net’ argument to tell the route program that this entry is a network route. Your other choice here is a ‘-host’ route which is a route that is specific to one IP address.

This route will enable you to establish IP connections with all of the hosts on your ethernet segment. But what about all of the IP hosts that aren’t on your ethernet segment ?

It would be a very difficult job to have to add routes to every possible destination network, so there is a special trick that is used to simplify this task. The trick is called the ‘default’ route. The default route matches every possible destination, but poorly, so that if any other entry exists that matches the required address it will be used instead of the default route. The idea of the default route is simply to enable you to say ”and everything else should go here”. In the example I’ve contrived you would use an entry like:

     # route add default gw 192.168.1.1 eth0

The ‘gw’ argument tells the route command that the next argument is the IP address, or name, of a gateway or router machine which all datagrams matching this entry should be directed to for further routing.

So, your complete configuration would look like:

     # ifconfig eth0 192.168.1.10 netmask 255.255.255.0 up
     # route add -net 192.168.1.0 netmask 255.255.255.0 eth0
     # route add default gw 192.168.1.1 eth0

If you take a close look at your network ‘rc’ files you will find that at least one of them looks very similar to this. This is a very common configuration.

Let’s now look at a slightly more complicated routing configuration. Let’s imagine we are configuring the router we looked at earlier, the one supporting the PPP link to the Internet and the lan segments feeding the workstations in the office. Lets imagine the router has three ethernet segments and one PPP link. Our routing configuration would look something like:

     # route add -net 192.168.1.0 netmask 255.255.255.0 eth0
     # route add -net 192.168.2.0 netmask 255.255.255.0 eth1
     # route add -net 192.168.3.0 netmask 255.255.255.0 eth2
     # route add default ppp0

Each of the workstations would use the simpler form presented above, only the router needs to specify each of the network routes seperately because for the workstations the default route mechanism will capture all of them letting the router worry about splitting them up appropriately. You may be wondering why the default route presented doesn’t specify a ‘gw’. The reason for this is simple, serial link protocols such as PPP and slip only ever have two hosts on their network, one at each end. To specify the host at the other end of the link as the gateway is pointless and redundant as there is no other choice, so you do not need to specify a gateway for these types of network connections. Other network types such as ethernet, arcnet or token ring do require the gateway to be specified as these networks support large numbers of hosts on them.

5.7.1 So what does the routed program do ?

The routing configuration described above is best suited to simple network arrangements where there are only ever single possible paths to destinations. When you have a more complex network arrangement things get a little more complicated. Fortunately for most of you this won’t be an issue.

The big problem with ‘manual routing’ or ‘static routing’ as described, is that if a machine or link fails in your network then the only way you can direct your datagrams another way, if another way exists, is by manually intervening and executing the appopriate commands. Naturally this is clumsy, slow, impractical and hazard prone. Various techniques have been developed to automatically adjust routing tables in the event of network failures where there are alternate routes, all of these techniques are loosely grouped by the term ‘dynamic routing protocols’.

You may have heard of some of the more common dynamic routing protocols. The most common are probably RIP (Routing Information Protocol) and OSPF (Open Shortest Path First Protocol). The Routing Information Protocol is very common on small networks such as small-medium sized corporate networks or building networks. OSPF is more modern and more capable at handling large network configurations and better suited to environments where there is a large number of possible paths through the network. Common implementations of these protocols are: ‘routed’ - RIP and ‘gated’ - RIP, OSPF and others. The ‘routed’ program is normally supplied with your Linux distribution or is included in the ‘NetKit’ package detailed above.

An example of where and how you might use a dynamic routing protocol might look something like the following:

         192.168.1.0 /                         192.168.2.0 /
            255.255.255.0                         255.255.255.0
          -                                     -
          |                                     |
          |   /-----\                 /-----\   |
          |   |     |ppp0   //    ppp0|     |   |
     eth0 |---|  A  |------//---------|  B  |---| eth0
          |   |     |     //          |     |   |
          |   \-----/                 \-----/   |
          |      \ ppp1             ppp1 /      |
          -       \                     /       -
                   \                   /
                    \                 /
                     \               /
                      \             /
                       \           /
                        \         /
                         \       /
                          \     /
                       ppp0\   /ppp1
                          /-----\
                          |     |
                          |  C  |
                          |     |
                          \-----/
                             |eth0
                             |
                        |---------|
                        192.168.3.0 /
                           255.255.255.0

We have three routers A, B and C. Each supports one ethernet segment with a Class C IP network (netmask 255.255.255.0). Each router also has a PPP link to each of the other routers. The network forms a triangle.

It should be clear that the routing table at router A could look like:

     # route add -net 192.168.1.0 netmask 255.255.255.0 eth0
     # route add -net 192.168.2.0 netmask 255.255.255.0 ppp0
     # route add -net 192.168.3.0 netmask 255.255.255.0 ppp1

This would work just fine until the link between router A and B should fail. If that link failed then with the routing entry shown above hosts on the ethernet segment of A could not reach hosts on the ethernet segment on B because their datagram would be directed to router A’s ppp0 link which is broken. They could still continue to talk to hosts on the ethernet segment of C and hosts on the C’s ethernet segment could still talk to hosts on B’s ethernet segment because the link between B and C is still intact.

But wait, if A can talk to C and C can still talk to B, why shouldn’t A route its datagrams for B via C and let C send them to B ? This is exactly the sort of problem that dynamic routing protocols like RIP were designed to solve. If each of the routers A, B and C were running a routing daemon then their routing tables would be automatically adjusted to reflect the new state of the network should any one of the links in the network fail. To configure such a network is simple, at each router you need only do two things. In this case for Router A:

     # route add -net 192.168.1.0 netmask 255.255.255.0 eth0
     # /usr/sbin/routed

The ‘routed’ routing daemon automatically finds all active network ports when it starts and sends and listens for messages on each of the network devices to allow it to determine and update the routing table on the host.

This has been a very brief explanation of dynamic routing and where you would use it. If you want more information then you should refer to the suggested references listed at the top of the document.

The important points relating to dynamic routing are:

1. You only need to run a dynamic routing protocol daemon when your Linux machine has the possibility of selecting multiple possible routes to a destination.

2. The dynamic routing daemon will automatically modify your routing table to adjust to changes in your network.

3. RIP is suited to small to medium sized networks.

5.8 Configuring your network servers and services.

Network servers and services are those programs that allow a remote user to make user of your Linux machine. Server programs listen on network ports. Network ports are a means of addressing a particular service on any particular host and are how a server knows the difference between an incoming telnet connection and an incoming ftp connection. The remote user establishes a network connection to your machine and the server program, the network daemon program, listening on that port accepts the connection and executes. There are two ways that network daemons may operate. Both are commonly employed in practice. The two ways are:

standalone the network daemon program listens on the designated network port and when an incoming connection is made it manages the network connection itself to provide the service.

slave to the inetd server the inetd server is a special network daemon program that specialises in managing incoming network connections. It has a configuration file which tells it what program needs to be run when an incoming connection is received. Any service port may be configured for either of the tcp or udp protcols. The ports are described in another file that we will talk about soon.

There are two important files that we need to configure. They are the /etc/services file which assigns names to port numbers and the /etc/inetd.conf file which is the configuration file for the inetd network daemon.

5.8.1 /etc/services

The /etc/services file is a simple database that associates a human friendly name to a machine friendly service port. Its format is quite simple. The file is a text file with each line representing and entry in the database. Each entry is comprised of three fields seperated by any number of whitespace (tab or space) characters. The fields are:

name      port/protocol        aliases     # comment

name a single word name that represents the service being described.

port/protocol this field is split into two subfields.

port a number that specifies the port number the named service will be available on. Most of the common services have assigned service numbers. These are described in RFC-1340.

protocol this subfield may be set to either tcp or udp.

It is important to note that an entry of 18/tcp is very different from an entry of 18/udp and that there is no technical reason why the same service needs to exist on both. Normally common sense prevails and it is only if a particular service is available via both tcp and udp that you will see an entry for both.

aliases other names that may be used to refer to this service entry.

Any text appearing in a line after a ‘#’ character is ignored and treated as a comment.

An example /etc/services file. All modern linux distributions provide a good /etc/services file. Just in case you happen to be building a machine from the ground up, here is a copy of the /etc/services file supplied with the Debian <http://www.debian.org/> distribution.

     # /etc/services:
     # $Id: services,v 1.3 1996/05/06 21:42:37 tobias Exp $
     #
     # Network services, Internet style
     #
     # Note that it is presently the policy of IANA to assign a single well-known
     # port number for both TCP and UDP; hence, most entries here have two entries
     # even if the protocol doesn't support UDP operations.
     # Updated from RFC 1340, ``Assigned Numbers'' (July 1992).  Not all ports
     # are included, only the more common ones.

5. Generic Network Configuration Information. 25

tcpmux          1/tcp                           # TCP port service multiplexer
echo            7/tcp
echo            7/udp
discard         9/tcp           sink null
discard         9/udp           sink null
systat          11/tcp          users
daytime         13/tcp
daytime         13/udp
netstat         15/tcp
qotd            17/tcp          quote
msp             18/tcp                          # message send protocol
msp             18/udp                          # message send protocol
chargen         19/tcp          ttytst source
chargen         19/udp          ttytst source
ftp-data        20/tcp
ftp             21/tcp
ssh             22/tcp                          # SSH Remote Login Protocol
ssh             22/udp                          # SSH Remote Login Protocol
telnet          23/tcp
# 24 - private
smtp            25/tcp          mail
# 26 - unassigned
time            37/tcp          timserver
time            37/udp          timserver
rlp             39/udp          resource        # resource location
nameserver      42/tcp          name            # IEN 116
whois           43/tcp          nicname
re-mail-ck      50/tcp                          # Remote Mail Checking Protocol
re-mail-ck      50/udp                          # Remote Mail Checking Protocol
domain          53/tcp          nameserver      # name-domain server
domain          53/udp          nameserver
mtp             57/tcp                          # deprecated
bootps          67/tcp                          # BOOTP server
bootps          67/udp
bootpc          68/tcp                          # BOOTP client
bootpc          68/udp
tftp            69/udp
gopher          70/tcp                          # Internet Gopher
gopher          70/udp
rje             77/tcp          netrjs
finger          79/tcp
www             80/tcp          http            # WorldWideWeb HTTP
www             80/udp                          # HyperText Transfer Protocol
link            87/tcp          ttylink
kerberos        88/tcp          kerberos5 krb5  # Kerberos v5
kerberos        88/udp          kerberos5 krb5  # Kerberos v5
supdup          95/tcp

# 100 - reserved
hostnames       101/tcp         hostname        # usually from sri-nic
iso-tsap        102/tcp         tsap            # part of ISODE.
csnet-ns        105/tcp         cso-ns          # also used by CSO name server
csnet-ns        105/udp         cso-ns
rtelnet         107/tcp                         # Remote Telnet
rtelnet         107/udp
pop-2           109/tcp         postoffice      # POP version 2
pop-2           109/udp
pop-3           110/tcp                         # POP version 3
pop-3           110/udp
sunrpc          111/tcp         portmapper      # RPC 4.0 portmapper TCP
sunrpc          111/udp         portmapper      # RPC 4.0 portmapper UDP
auth            113/tcp         authentication tap ident
sftp            115/tcp
uucp-path       117/tcp
nntp            119/tcp         readnews untp   # USENET News Transfer Protocol
ntp             123/tcp
ntp             123/udp                         # Network Time Protocol
netbios-ns      137/tcp                         # NETBIOS Name Service
netbios-ns      137/udp
netbios-dgm     138/tcp                         # NETBIOS Datagram Service
netbios-dgm     138/udp
netbios-ssn     139/tcp                         # NETBIOS session service
netbios-ssn     139/udp
imap2           143/tcp                         # Interim Mail Access Proto v2
imap2           143/udp
snmp            161/udp                         # Simple Net Mgmt Proto
snmp-trap       162/udp         snmptrap        # Traps for SNMP
cmip-man        163/tcp                         # ISO mgmt over IP (CMOT)
cmip-man        163/udp
cmip-agent      164/tcp
cmip-agent      164/udp
xdmcp           177/tcp                         # X Display Mgr. Control Proto
xdmcp           177/udp
nextstep        178/tcp         NeXTStep NextStep       # NeXTStep window
nextstep        178/udp         NeXTStep NextStep       # server
bgp             179/tcp                         # Border Gateway Proto.
bgp             179/udp
prospero        191/tcp                         # Cliff Neuman's Prospero
prospero        191/udp
irc             194/tcp                         # Internet Relay Chat
irc             194/udp
smux            199/tcp                         # SNMP Unix Multiplexer
smux            199/udp
at-rtmp         201/tcp                         # AppleTalk routing
at-rtmp         201/udp
at-nbp          202/tcp                         # AppleTalk name binding

at-nbp          202/udp
at-echo         204/tcp                         # AppleTalk echo
at-echo         204/udp
at-zis          206/tcp                         # AppleTalk zone information
at-zis          206/udp
z3950           210/tcp         wais            # NISO Z39.50 database
z3950           210/udp         wais
ipx             213/tcp                         # IPX
ipx             213/udp
imap3           220/tcp                         # Interactive Mail Access
imap3           220/udp                         # Protocol v3
ulistserv       372/tcp                         # UNIX Listserv
ulistserv       372/udp
#
# UNIX specific services
#
exec            512/tcp
biff            512/udp         comsat
login           513/tcp
who             513/udp         whod
shell           514/tcp         cmd             # no passwords used
syslog          514/udp
printer         515/tcp         spooler         # line printer spooler
talk            517/udp
ntalk           518/udp
route           520/udp         router routed   # RIP
timed           525/udp         timeserver
tempo           526/tcp         newdate
courier         530/tcp         rpc
conference      531/tcp         chat
netnews         532/tcp         readnews
netwall         533/udp                         # -for emergency broadcasts
uucp            540/tcp         uucpd           # uucp daemon
remotefs        556/tcp         rfs_server rfs  # Brunhoff remote filesystem
klogin          543/tcp                         # Kerberized `rlogin' (v5)
kshell          544/tcp         krcmd           # Kerberized `rsh' (v5)
kerberos-adm    749/tcp                         # Kerberos `kadmin' (v5)
#
webster         765/tcp                         # Network dictionary
webster         765/udp
#
# From ``Assigned Numbers'':
#
#> The Registered Ports are not controlled by the IANA and on most systems
#> can be used by ordinary user processes or programs executed by ordinary
#> users.
#
#> Ports are used in the TCP [45,106] to name the ends of logical

#> connections which carry long term conversations.  For the purpose of
#> providing services to unknown callers, a service contact port is
#> defined.  This list specifies the port used by the server process as its
#> contact port.  While the IANA can not control uses of these ports it
#> does register or list uses of these ports as a convienence to the
#> community.
#
ingreslock      1524/tcp
ingreslock      1524/udp
prospero-np     1525/tcp                # Prospero non-privileged
prospero-np     1525/udp
rfe             5002/tcp                # Radio Free Ethernet
rfe             5002/udp                # Actually uses UDP only
bbs             7000/tcp                # BBS service
#
#
# Kerberos (Project Athena/MIT) services
# Note that these are for Kerberos v4 and are unofficial.  Sites running
# v4 should uncomment these and comment out the v5 entries above.
#
kerberos4       750/udp         kdc     # Kerberos (server) udp
kerberos4       750/tcp         kdc     # Kerberos (server) tcp
kerberos_master 751/udp                 # Kerberos authentication
kerberos_master 751/tcp                 # Kerberos authentication
passwd_server   752/udp                 # Kerberos passwd server
krb_prop        754/tcp                 # Kerberos slave propagation
krbupdate       760/tcp         kreg    # Kerberos registration
kpasswd         761/tcp         kpwd    # Kerberos "passwd"
kpop            1109/tcp                # Pop with Kerberos
knetd           2053/tcp                # Kerberos de-multiplexor
zephyr-srv      2102/udp                # Zephyr server
zephyr-clt      2103/udp                # Zephyr serv-hm connection
zephyr-hm       2104/udp                # Zephyr hostmanager
eklogin         2105/tcp                # Kerberos encrypted rlogin
#
# Unofficial but necessary (for NetBSD) services
#
supfilesrv      871/tcp                 # SUP server
supfiledbg      1127/tcp                # SUP debugging
#
# Datagram Delivery Protocol services
#
rtmp            1/ddp                   # Routing Table Maintenance Protocol
nbp             2/ddp                   # Name Binding Protocol
echo            4/ddp                   # AppleTalk Echo Protocol
zip             6/ddp                   # Zone Information Protocol
#
# Debian GNU/Linux services

rmtcfg          1236/tcp                # Gracilis Packeten remote config server
xtel            1313/tcp                # french minitel
cfinger         2003/tcp                # GNU Finger
postgres        4321/tcp                # POSTGRES
mandelspawn     9359/udp        mandelbrot      # network mandelbrot
# Local services
5.8.2 /etc/inetd.conf

The /etc/inetd.conf file is the configuration file for the inetd server daemon. Its function is to tell inetd what to do when it receives a connection request for a particular service. For each service that you wish to accept connections for you must tell inetd what network server daemon to run and how to run it.

Its format is also fairly simple. It is a text file with each line describing a service that you wish to provide. Any text in a line following a ‘#’ is ignored and considered a comment. Each line contains seven fields seperated by any number of whitespace (tab or space) characters. The general format is as follows:

service  socket_type  proto  flags  user  server_path  server_args

service is the service relevant to this configuration as taken from the /etc/services file.

socket type this field describes the type of socket that this entry will consider relevant, allowable values are: stream, dgram, raw, rdm, or seqpacket. This is a little technical in nature, but as a rule of thumb nearly all tcp based services use stream and nearly all udp based services use dgram. It is only very special types of server daemons that would use any of the other values.

proto the protocol to considered valid for this entry. This should match the appropriate entry in the /etc/services file and will typically be either tcp or udp. Sun RPC (Remote Procedure Call) based servers will use rpc/tcp or rpc/udp.

flags there are really only two possible settings for this field. This field setting tells inetd whether the network server program frees the socket after it has been started and therefore whether inetd can start another one on the next connection request, or whether inetd should wait and assume that any server daemon already running will handle the new connection request. Again this is a little tricky to work out, but as a rule of thumb all tcp servers should have this entry set to nowait and most udp servers should have this entry set to wait. Be warned there are some notable exceptions to this, so let the example guide you if you are not sure.

user this field describes which user account from /etc/passwd will be set as the owner of the network daemon when it is started. This is often useful if you want to safeguard against security risks. You can set the user of an entry to the nobody user so that if the network server security is breached the possible damage is minimised. Typically this field is set to root though, because many servers require root priveledges in order to function correctly.

server path

this field is pathname to the actual server program to execute for this entry.

server args

this field comprises the rest of the line and is optional. This field is where you place any command line arguments that you wish to pass to the server daemon program when it is launched.

An example /etc/inetd.conf As for the /etc/services file all modern distributions will include a good /etc/inetd.conf file for you to work with. Here, for completeness is the /etc/inetd.conf file from

the Debian <http://www.debian.org/> distribution.
# /etc/inetd.conf:  see inetd(8) for further informations.
#
# Internet server configuration database
#
#
# Modified for Debian by Peter Tobias <tobias@et-inf.fho-emden.de>
#
# <service_name> <sock_type> <proto> <flags> <user> <server_path> <args>
#
# Internal services
#
#echo           stream  tcp     nowait  root    internal
#echo           dgram   udp     wait    root    internal
discard         stream  tcp     nowait  root    internal
discard         dgram   udp     wait    root    internal
daytime         stream  tcp     nowait  root    internal
daytime         dgram   udp     wait    root    internal
#chargen        stream  tcp     nowait  root    internal
#chargen        dgram   udp     wait    root    internal
time            stream  tcp     nowait  root    internal
time            dgram   udp     wait    root    internal
#
# These are standard services.
#
telnet  stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.telnetd
ftp     stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.ftpd
#fsp    dgram   udp     wait    root    /usr/sbin/tcpd  /usr/sbin/in.fspd
#
# Shell, login, exec and talk are BSD protocols.
#
shell   stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.rshd
login   stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.rlogind

#exec   stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.rexecd
talk    dgram   udp     wait    root    /usr/sbin/tcpd  /usr/sbin/in.talkd
ntalk   dgram   udp     wait    root    /usr/sbin/tcpd  /usr/sbin/in.ntalkd
#
# Mail, news and uucp services.
#
smtp    stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.smtpd
#nntp   stream  tcp     nowait  news    /usr/sbin/tcpd  /usr/sbin/in.nntpd
#uucp   stream  tcp     nowait  uucp    /usr/sbin/tcpd  /usr/lib/uucp/uucico
#comsat dgram   udp     wait    root    /usr/sbin/tcpd  /usr/sbin/in.comsat
#
# Pop et al
#
#pop-2  stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.pop2d
#pop-3  stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.pop3d
#
# `cfinger' is for the GNU finger server available for Debian.  (NOTE: The
# current implementation of the `finger' daemon allows it to be run as `root'.)
#
#cfinger stream tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.cfingerd
#finger stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.fingerd
#netstat        stream  tcp     nowait  nobody  /usr/sbin/tcpd  /bin/netstat
#systat stream  tcp     nowait  nobody  /usr/sbin/tcpd  /bin/ps -auwwx
#
# Tftp service is provided primarily for booting.  Most sites
# run this only on machines acting as "boot servers."
#
#tftp   dgram   udp     wait    nobody  /usr/sbin/tcpd  /usr/sbin/in.tftpd
#tftp   dgram   udp     wait    nobody  /usr/sbin/tcpd  /usr/sbin/in.tftpd /boot
#bootps dgram   udp     wait    root    /usr/sbin/bootpd        bootpd -i -t 120
#
# Kerberos authenticated services (these probably need to be corrected)
#
#klogin         stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.rlogind -k
#eklogin        stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.rlogind -k -x
#kshell         stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/in.rshd -k
#
# Services run ONLY on the Kerberos server (these probably need to be corrected)
#
#krbupdate      stream tcp      nowait  root    /usr/sbin/tcpd  /usr/sbin/registerd
#kpasswd        stream  tcp     nowait  root    /usr/sbin/tcpd  /usr/sbin/kpasswdd
#
# RPC based services
#
#mountd/1       dgram   rpc/udp wait    root    /usr/sbin/tcpd  /usr/sbin/rpc.mountd
#rstatd/1-3     dgram   rpc/udp wait    root    /usr/sbin/tcpd  /usr/sbin/rpc.rstatd
#rusersd/2-3    dgram   rpc/udp wait    root    /usr/sbin/tcpd  /usr/sbin/rpc.rusersd
#walld/1        dgram   rpc/udp wait    root    /usr/sbin/tcpd  /usr/sbin/rpc.rwalld

#
# End of inetd.conf.
ident           stream  tcp     nowait  nobody  /usr/sbin/identd        identd -i

5.9 Other miscellaneous network related configuration files.

There are a number of miscellaneous files relating to network configuration under linux that you might be interested in. You may never have to modify these files, but it is worth describing them so you know what they contain and what they are for.

5.9.1 /etc/protocols

The /etc/protocols file is a database that maps protocol id numbers against protocol names. This is used by programmers to allow them to specify protocols by name in their programs and also by some programs such as tcpdump to allow them to display names instead of numbers in their output. The general syntax of the file is:

protocolname  number  aliases

The /etc/protocols file supplied with the Debian <http://www.debian.org/> distribution is as follows:

# /etc/protocols:
# $Id: protocols,v 1.1 1995/02/24 01:09:41 imurdock Exp $
#
# Internet (IP) protocols
#
#       from: @(#)protocols     5.1 (Berkeley) 4/17/89
#
# Updated for NetBSD based on RFC 1340, Assigned Numbers (July 1992).
ip      0       IP              # internet protocol, pseudo protocol number
icmp    1       ICMP            # internet control message protocol
igmp    2       IGMP            # Internet Group Management
ggp     3       GGP             # gateway-gateway protocol
ipencap 4       IP-ENCAP        # IP encapsulated in IP (officially ``IP'')
st      5       ST              # ST datagram mode
tcp     6       TCP             # transmission control protocol
egp     8       EGP             # exterior gateway protocol
pup     12      PUP             # PARC universal packet protocol
udp     17      UDP             # user datagram protocol
hmp     20      HMP             # host monitoring protocol
xns-idp 22      XNS-IDP         # Xerox NS IDP
rdp     27      RDP             # "reliable datagram" protocol
iso-tp4 29      ISO-TP4         # ISO Transport Protocol class 4
xtp     36      XTP             # Xpress Tranfer Protocol
ddp     37      DDP             # Datagram Delivery Protocol

     idpr-cmtp       39      IDPR-CMTP       # IDPR Control Message Transport
     rspf    73      RSPF            # Radio Shortest Path First.
     vmtp    81      VMTP            # Versatile Message Transport
     ospf    89      OSPFIGP         # Open Shortest Path First IGP
     ipip    94      IPIP            # Yet Another IP encapsulation
     encap   98      ENCAP           # Yet Another IP encapsulation
5.9.2 /etc/networks

The /etc/networks file has a similar function to that of the /etc/hosts file. It provides a simple database of network names against network addresses. Its format differs in that there may be only two fields per line and that the fields are coded as:

     networkname networkaddress

An example might look like:

     loopnet    127.0.0.0
     localnet   192.168.0.0
     amprnet    44.0.0.0

When you use commands like the route command, if a destination is a network and that network has an entry in the /etc/networks file then the route command will display that network name instead of its address.

5.10 Network Security and access control.

Let me start this section by warning you that securing your machine and network against malicious attack is a complex art. I do not consider myself an expert in this field at all and while the following mechanisms I describe will help, if you are serious about security then I recommend you do some research of your own into the subject. There are many good references on the Internet relating to the subject.

An important rule of thumb is: ‘Don’t run servers you don’t intend to use’. Many distributions come configured with all sorts of services configured and automatically started. To ensure even a minimum level of safety you should go through your /etc/inetd.conf file and comment out (place a ‘#’ at the start of the line) any entries for services you don’t intend to use. Good candidates are services such as: shell, login, exec, uucp, ftp and informational services such as finger, netstat and systat.

There are all sorts of security and access control mechanisms, I’ll describe the most elementary of them.

5.10.1 /etc/ftpusers

The /etc/ftpusers file is a simple mechanism that allows you to deny certain users from logging into your machine via ftp. The /etc/ftpusers file is read by the ftp daemon program (ftpd) when an incoming ftp connection is received. The file is a simple list of those users who are disallowed from logging in. It might looks something like:

     # /etc/ftpusers - users not allowed to login via ftp
     root
     uucp
     bin
     mail

5.10.2 /etc/securetty

The /etc/securetty file allows you to specify which tty devices root is allowed to login on. The /etc/securetty file is read by the login program (usually /bin/login). Its format is a list of the tty devices names allowed, on all others root login is disallowed:

     # /etc/securetty - tty's on which root is allowed to login
     tty1
     tty2
     tty3
     tty4

5.10.3 The tcpd hosts access control mechanism.

The tcpd program you will have seen listed in the same /etc/inetd.conf provides logging and access control mechanisms to services it is configured to protect.

When it is invoked by the inetd program it reads two files containing access rules and either allows or denies access to the server it is protecting accordingly.

It will search the rules files until the first match is found. If no match is found then it assumes that access should be allowed to anyone. The files it searches in sequence are: /etc/hosts.allow, /etc/hosts.deny. I’ll describe each of these in turn. For a complete description of this facility you should refer to the appopriate man pages (hosts access(5) is a good starting point).

/etc/hosts.allow The /etc/hosts.allow file is a configuration file of the /usr/sbin/tcpd program. The hosts.allow file contains rules describing which hosts are allowed access to a service on your machine.

The file format is quite simple:

     # /etc/hosts.allow
     #
     # <service list>: <host list> [: command]
service list

is a comma delimited list of server names that this rule applies to. Example server names are: ftpd,

     telnetd and fingerd.
host list

is a comma delimited list of host names. You may also use IP addresses here. You may additionally specify hostnames or addresses using wildcard characters to match groups of hosts. Examples include: gw.vk2ktj.ampr.org to match a specific host, .uts.edu.au to match any hostname ending in that string, 44. to match any IP address commencing with those digits. There are some special tokens to simplify configuration, some of these are: ALL matches every host, LOCAL matches any host whose name does not contain a ‘.’ ie is in the same domain as your machine and PARANOID matches any host whose name does not match its address (name spoofing). There is one last token that is also useful. The EXCEPT token allows you to provide a list with exceptions. This will be covered in an example later.

command

is an optional parameter. This parameter is the full pathname of a command that would be executed everytime this rule is matched. It could for example run a command that would attempt to identify who is logged onto the connecting host, or to generate a mail message or some other warning to a system administrator that someone is attempting to connect. There are a number of expansions that may be included, some common examples are: %h expands to the name of the connecting host or address if it doesn’t have a name, %d the daemon name being called.

An example:

# /etc/hosts.allow
#
# Allow mail to anyone
in.smtpd: ALL
# All telnet and ftp to only hosts within my domain and my host at home.
telnetd, ftpd: LOCAL, myhost.athome.org.au
# Allow finger to anyone but keep a record of who they are.
fingerd: ALL: (finger @%h | mail -s "finger from %h" root)

/etc/hosts.deny The /etc/hosts.deny file is a configuration file of the /usr/sbin/tcpd program. The hosts.deny file contains rules describing which hosts are disallowed access to a service on your machine.

A simple sample would look something like this:

# /etc/hosts.deny
#
# Disallow all hosts with suspect hostnames
ALL: PARANOID
#
# Disallow all hosts.
ALL: ALL

The PARANOID entry is really redundant because the other entry traps everything in any case. Either of these entry would make a reasonable default depending on your particular requirement.

Having an ALL: ALL default in the /etc/hosts.deny and then specifically enabling on those services and hosts that you want in the /etc/hosts.allow file is the safest configuration.

5.10.4 /etc/hosts.equiv

The hosts.equiv file is used to grant certain hosts and users access rights to accounts on your machine without having to supply a password. This is useful in a secure environment where you control all machines, but is a security hazard otherwise. Your machine is only as secure as the least secure of the trusted hosts. To maximise security, don’t use this mechanism and encourage your users not to use the .rhosts file as well.

5.10.5 Configure your ftp daemon properly.

Many sites will be interested in running an anonymous ftp server to allow other people to upload and download files without requiring a specific userid. If you decide to offer this facility make sure you configure the ftp daemon properly for anonymous access. Most man pages for ftpd(8) describe in some length how to go about this. You should always ensure that you follow these instructions. An important tip is to not use a copy of your /etc/passwd file in the anonymous account /etc directory, make sure you strip out all account details except those that you must have, otherwise you will be vulnerable to brute force password cracking techniques.

5.10.6 Network Firewalling.

Not allowing datagrams to even reach your machine or servers is an excellent means of security. This is covered in depth in the Firewall-HOWTO <Firewall-HOWTO.html>.

5.10.7 Other suggestions.

Here are some other, potentially religious suggestions for you to consider.

sendmail despite its popularity the sendmail daemon appears with frightening regularity on security warning announcements. Its up to you, but I choose not to run it.

NFS and other Sun RPC services be wary of these. There are all sorts of possible exploits for these services. It is difficult finding an option to services like NFS, but if you configure them, make sure you are careful with who you allow mount rights to.

6 Network Technology Specific Information.

The following subsections are specific to particular network technologies. The information contained in these sections does not necessarily apply to any other type of network technology.

6.1 ARCNet

ARCNet device names are ‘arc0e’, ‘arc1e’, ‘arc2e’ etc. or ‘arc0s’, ‘arc1s’, ‘arc2s’ etc. The first card detected by the kernel is assigned ‘arc0e’ or ‘arc0s’ and the rest are assigned sequentially in the order they are detected. The letter at the end signifies whether you’ve selected ethernet encapsulation packet format or RFC1051 packet format.

Kernel Compile Options:

     Network device support  --->
         [*] Network device support
         <*> ARCnet support
         [ ]   Enable arc0e (ARCnet "Ether-Encap" packet format)
         [ ]   Enable arc0s (ARCnet RFC1051 packet format)

Once you have your kernel properly built to support your ethernet card then configuration of the card is easy.

Typically you would use something like:

     # ifconfig arc0e 192.168.0.1 netmask 255.255.255.0 up
     # route add -net 192.168.0.0 netmask 255.255.255.0 arc0e
Please refer to the /usr/src/linux/Documentation/networking/arcnet.txt and
/usr/src/linux/Documentation/networking/arcnet-hardware.txt files for further information.

ARCNet support was developed by Avery Pennarun, apenwarr@foxnet.net.

6.2 Appletalk (AF APPLETALK)

The Appletalk support has no special device names as it uses existing network devices.

Kernel Compile Options:

     Networking options  --->
         <*> Appletalk DDP

Appletalk support allows your Linux machine to interwork with Apple networks. An important use for this is to share resources such as printers and disks between both your Linux and Apple computers. Additional software is required, this is called netatalk. Wesley Craig netatalk@umich.edu represents a team called the ‘Research Systems Unix Group’ at the University of Michigan and they have produced the netatalk package which provides software that implements the Appletalk protocol stack and some useful utilities. The netatalk package will either have been supplied with your Linux distribution, or you will have to ftp it from

its home site at the University of Michigan <ftp://terminator.rs.itd.umich.edu/unix/netatalk/>

To build and install the package do something like:

     # cd /usr/src
     # tar xvfz .../netatalk-1.4b2.tar.Z

     - You may want to edit the `Makefile' at this point, specifically to change
       the DESTDIR variable which defines where the files will be installed later.
       The default of /usr/local/atalk is fairly safe.
     # make
     - as root:
     # make install

6.2.1 Configuring the Appletalk software.

The first thing you need to do to make it all work is to ensure that the appropriate entries in the /etc/services file are present. The entries you need are:

     rtmp    1/ddp   # Routing Table Maintenance Protocol
     nbp     2/ddp   # Name Binding Protocol
     echo    4/ddp   # AppleTalk Echo Protocol
     zip     6/ddp   # Zone Information Protocol

The next step is to create the Appletalk configuration files in the /usr/local/atalk/etc directory (or wherever you installed the package).

The first file to create is the /usr/local/atalk/etc/atalkd.conf file. Initially this file needs only one line that gives the name of the network device that supports the network that your Apple machines are on:

     eth0

The Appletalk daemon program will add extra details after it is run.

6.2.2 Exporting a Linux filesystems via Appletalk.

You can export filesystems from your linux machine to the network so that Apple machine on the network can share them.

To do this you must configure the /usr/local/atalk/etc/AppleVolumes.system file. There is another configuration file called /usr/local/atalk/etc/AppleVolumes.default which has exactly the same format and describes which filesystems users connecting with guest priveledges will receive.

Full details on how to configure these files and what the various options are can be found in the afpd man page.

A simple example might look like:

     /tmp Scratch
     /home/ftp/pub "Public Area"

Which would export your /tmp filesystem as AppleShare Volume ‘Scratch’ and your ftp public directory as AppleShare Volume ‘Public Area’. The volume names are not mandatory, the daemon will choose some for you, but it won’t hurt to specify them anyway.

6.2.3 Sharing your Linux printer across Appletalk.

You can share your linux printer with your Apple machines quite simply. You need to run the papd program which is the Appletalk Printer Access Protocol Daemon. When you run this program it will accept requests from your Apple machines and spool the print job to your local line printer daemon for printing.

You need to edit the /usr/local/atalk/etc/papd.conf file to configure the daemon. The syntax of this file is the same as that of your usual /etc/printcap file. The name you give to the definition is registered with the Appletalk naming protocol, NBP.

A sample configuration might look like:

     TricWriter:\
        :pr=lp:op=cg:

Which would make a printer named ‘TricWriter’ available to your Appletalk network and all accepted jobs would be printed to the linux printer ‘lp’ (as defined in the /etc/printcap file) using lpd. The entry ‘op=cg’ says that the linux user ‘cg’ is the operator of the printer.

6.2.4 Starting the appletalk software.

Ok, you should now be ready to test this basic configuration. There is an rc.atalk file supplied with the netatalk package that should work ok for you, so all you should have to do is:

     # /usr/local/atalk/etc/rc.atalk

and all should startup and run ok. You should see no error messages and the software will send messages to the console indicating each stage as it starts.

6.2.5 Testing the appletalk software.

To test that the software is functioning properly, go to one of your Apple machines, pull down the Apple menu, select the Chooser, click on AppleShare, and your Linux box should appear.

6.2.6 Caveats of the appletalk software.

• You may need to start the Appletalk support before you configure your IP network. If you have problems starting the Appletalk programs, or if after you start them you have trouble with your IP network, then try starting the Appletalk software before you run your /etc/rc.d/rc.inet1 file.

• The afpd (Apple Filing Protocol Daemon) severely messes up your hard disk. Below the mount points it creates a couple of directories called .AppleDesktop and Network Trash Folder. Then, for each directory you access it will create a .AppleDouble below it so it can store resource forks, etc. So think twice before exporting /, you will have a great time cleaning up afterwards.

• The afpd program expects clear text passwords from the Macs. Security could be a problem, so be very careful when you run this daemon on a machine connected to the Internet, you have yourself to blame if somebody nasty does something bad.

• The existing diagnostic tools such as netstat and ifconfig don’t support Appletalk. The raw information is available in the /proc/net/ directory if you need it.

6.2.7 More information

For a much more detailed description of how to configure Appletalk for Linux refer to Anders Brownworth

Linux Netatalk-HOWTO page at thehamptons.com <http://thehamptons.com/anders/netatalk/>.

6.3 ATM

Werner Almesberger <werner.almesberger@lrc.di.epfl.ch> is managing a project to provide Asynchronous Transfer Mode support for Linux. Current information on the status of the project may be

obtained from: lrcwww.epfl.ch <http://lrcwww.epfl.ch/linux-atm/>.

6.4 AX25 (AF AX25)

AX.25 device names are ‘sl0’, ‘sl1’, etc. in 2.0.* kernels or ‘ax0’, ‘ax1’, etc. in 2.1.* kernels.

Kernel Compile Options:

     Networking options  --->
         [*] Amateur Radio AX.25 Level 2

The AX25, Netrom and Rose protocols are covered by the AX25-HOWTO <AX25-HOWTO.html>. These protocols are used by Amateur Radio Operators world wide in packet radio experimentation.

Most of the work for implementation of these protocols has been done by Jonathon Naylor,

jsn@cs.nott.ac.uk.

6.5 DECNet

Support for DECNet is currently being worked on. You should expect it to appear in a late 2.1.* kernel.

6.6 EQL - multiple line traffic equaliser

The EQL device name is ‘eql’. With the standard kernel source you may have only one EQL device per machine. EQL provides a means of utilising multiple point to point lines such as PPP, slip or plip as a single logical link to carry tcp/ip. Often it is cheaper to use multiple lower speed lines than to have one high speed line installed.

Kernel Compile Options:

     Network device support  --->
         [*] Network device support
         <*> EQL (serial line load balancing) support

To support this mechanism the machine at the other end of the lines must also support EQL. Linux, Livingstone Portmasters and newer dial-in servers support compatible facilities.

To configure EQL you will need the eql tools which are available from: sunsite.unc.edu

<ftp://sunsite.unc.edu/pub/linux/system/Serial/eql-1.2.tar.gz>.

Configuration is fairly straightforward. You start by configuring the eql interface. The eql interface is just like any other network device. You configure the IP address and mtu using the ifconfig utility, so something like:

     ifconfig eql 192.168.10.1 mtu 1006

Next you need to manually initiate each of the lines you will use. These may be any combination of point to point network devices. How you initiate the connections will depend on what sort of link they are, refer to the appropriate sections for further information.

Lastly you need to associate the serial link with the EQL device, this is called ‘enslaving’ and is done with the eql enslave command as shown:

     eql_enslave eql sl0 28800
     eql_enslave eql ppp0 14400

The ‘estimated speed’ parameter you supply eql enslave doesn’t do anything directly. It is used by the EQL driver to determine what share of the datagrams that device should receive, so you can fine tune the balancing of the lines by playing with this value.

To disassociate a line from an EQL device you use the eql emancipate command as shown:

     eql_emancipate eql sl0

You add routing as you would for any other point to point link, except your routes should refer to the eql device rather than the actual serial devices themselves, typically you would use:

     route add default eql

The EQL driver was developed by Simon Janes, simon@ncm.com.

6.7 Ethernet

Ethernet device names are ‘eth0’, ‘eth1’, ‘eth2’ etc. The first card detected by the kernel is assigned ‘eth0’ and the rest are assigned sequentially in the order they are detected.

To learn how to make your ethernet card working under Linux you should refer to the Ethernet-HOWTO

<Ethernet-HOWTO.html>.

Once you have your kernel properly built to support your ethernet card then configuration of the card is easy.

Typically you would use something like:

     # ifconfig eth0 192.168.0.1 netmask 255.255.255.0 up
     # route add -net 192.168.0.0 netmask 255.255.255.0 eth0

Most of the ethernet drivers were developed by Donald Becker, becker@CESDIS.gsfc.nasa.gov.

6.8 FDDI

FDDI device names are ‘fddi0’, ‘fddi1’, ‘fddi2’ etc. The first card detected by the kernel is assigned ‘fddi0’ and the rest are assigned sequentially in the order they are detected.

Lawrence V. Stefani, larry stefani@us.newbridge.com, has developed a driver for the Digital Equipment Corporation FDDI EISA and PCI cards.

Kernel Compile Options:

     Network device support  --->
         [*] FDDI driver support
         [*] Digital DEFEA and DEFPA adapter support

When you have your kernel built to support the FDDI driver and installed, configuration of the FDDI interface is almost identical that for an ethernet interface. You just specify the appropraite FDDI interface name in the ifconfig and route comands.

6.9 Frame Relay

The Frame Relay device names are ‘dlci00’, ‘dlci01’ etc for the DLCI encapsulation devices and ‘sdla0’, ‘sdla1’ etc for the FRAD(s).

Frame Relay is a new networking technology that is designed to suit data communications traffic that is of a ‘bursty’ or intermittent nature. You connect to a Frame Relay network using a Frame Relay Access Device (FRAD). The Linux Frame Relay supports IP over Frame Relay as described in RFC-1490.

Kernel Compile Options:

     Network device support  --->
         <*> Frame relay DLCI support (EXPERIMENTAL)
         (24)   Max open DLCI
         (8)   Max DLCI per device
         <*>   SDLA (Sangoma S502/S508) support

Mike McLagan, mike.mclagan@linux.org, developed the Frame Relay support and configuration tools.

Currently the only FRAD supported are the Sangoma Technologies <http://www.sangoma.com/> S502A,

S502E and S508.

6. Network Technology Specific Information. 43

To configure the FRAD and DLCI devices after you have rebuilt your kernel you will need the Frame Relay configuration tools. These are available from ftp.invlogic.com

<ftp://ftp.invlogic.com/pub/linux/fr/frad-0.15.tgz>. Compiling and installing the tools is

straightforward, but the lack of a top level Makefile makes it a fairly manual process:

# cd /usr/src
# tar xvfz .../frad-0.15.tgz
# cd frad-0.15
# for i in common dlci frad; make -C $i clean; make -C $i; done
# mkdir /etc/frad
# install -m 644 -o root -g root bin/*.sfm /etc/frad
# install -m 700 -o root -g root frad/fradcfg /sbin
# install -m 700 -o root -g root dlci/dlcicfg /sbin

After installing the tools you need to create an /etc/frad/router.conf file. You can use this template, which is a modified version of one of the example files:

# /etc/frad/router.conf
# This is a template configuration for frame relay.
# All tags are included. The default values are based on the code
# supplied with the DOS drivers for the Sangoma S502A card.
#
# A '#' anywhere in a line constitutes a comment
# Blanks are ignored (you can indent with tabs too)
# Unknown [] entries and unknown keys are ignored
#
[Devices]
Count=1                 # number of devices to configure
Dev_1=sdla0             # the name of a device
#Dev_2=sdla1            # the name of a device
# Specified here, these are applied to all devices and can be overriden for
# each individual board.
#
Access=CPE
Clock=Internal
KBaud=64
Flags=TX
#
# MTU=1500              # Maximum transmit IFrame length, default is 4096
# T391=10               # T391 value    5 - 30, default is 10
# T392=15               # T392 value    5 - 30, default is 15
# N391=6                # N391 value    1 - 255, default is 6
# N392=3                # N392 value    1 - 10, default is 3
# N393=4                # N393 value    1 - 10, default is 4
# Specified here, these set the defaults for all boards

# CIRfwd=16             # CIR forward   1 - 64
# Bc_fwd=16             # Bc forward    1 - 512
# Be_fwd=0              # Be forward    0 - 511
# CIRbak=16             # CIR backward  1 - 64
# Bc_bak=16             # Bc backward   1 - 512
# Be_bak=0              # Be backward   0 - 511

#
#
# Device specific configuration
#
#
#
# The first device is a Sangoma S502E
#
[sdla0]
Type=Sangoma            # Type of the device to configure, currently only
                        # SANGOMA is recognised
#
# These keys are specific to the 'Sangoma' type
#
# The type of Sangoma board - S502A, S502E, S508
Board=S502E
#
# The name of the test firmware for the Sangoma board
# Testware=/usr/src/frad-0.10/bin/sdla_tst.502
#
# The name of the FR firmware
# Firmware=/usr/src/frad-0.10/bin/frm_rel.502
#
Port=360                # Port for this particular card
Mem=C8                  # Address of memory window, A0-EE, depending on card
IRQ=5                   # IRQ number, do not supply for S502A
DLCIs=1                 # Number of DLCI's attached to this device
DLCI_1=16               # DLCI #1's number, 16 - 991
# DLCI_2=17
# DLCI_3=18
# DLCI_4=19
# DLCI_5=20
#
# Specified here, these apply to this device only,
# and override defaults from above
#
# Access=CPE            # CPE or NODE, default is CPE
# Flags=TXIgnore,RXIgnore,BufferFrames,DropAborted,Stats,MCI,AutoDLCI
# Clock=Internal        # External or Internal, default is Internal

# Baud=128              # Specified baud rate of attached CSU/DSU
# MTU=2048              # Maximum transmit IFrame length, default is 4096
# T391=10               # T391 value    5 - 30, default is 10
# T392=15               # T392 value    5 - 30, default is 15
# N391=6                # N391 value    1 - 255, default is 6
# N392=3                # N392 value    1 - 10, default is 3
# N393=4                # N393 value    1 - 10, default is 4
#
# The second device is some other card
#
# [sdla1]
# Type=FancyCard        # Type of the device to configure.
# Board=                # Type of Sangoma board
# Key=Value             # values specific to this type of device

#
# DLCI Default configuration parameters
# These may be overridden in the DLCI specific configurations
#
CIRfwd=64               # CIR forward   1 - 64
# Bc_fwd=16             # Bc forward    1 - 512
# Be_fwd=0              # Be forward    0 - 511
# CIRbak=16             # CIR backward  1 - 64
# Bc_bak=16             # Bc backward   1 - 512
# Be_bak=0              # Be backward   0 - 511
#
# DLCI Configuration
# These are all optional. The naming convention is
# [DLCI_D<devicenum>_<DLCI_Num>]
#
[DLCI_D1_16]
# IP=
# Net=
# Mask=
# Flags defined by Sangoma: TXIgnore,RXIgnore,BufferFrames
# DLCIFlags=TXIgnore,RXIgnore,BufferFrames
# CIRfwd=64
# Bc_fwd=512
# Be_fwd=0
# CIRbak=64
# Bc_bak=512
# Be_bak=0
[DLCI_D2_16]

     # IP=
     # Net=
     # Mask=
     # Flags defined by Sangoma: TXIgnore,RXIgnore,BufferFrames
     # DLCIFlags=TXIgnore,RXIgnore,BufferFrames
     # CIRfwd=16
     # Bc_fwd=16
     # Be_fwd=0
     # CIRbak=16
     # Bc_bak=16
     # Be_bak=0

When you’ve built your /etc/frad/router.conf file the only step remaining is to configure the actual devices themselves. This is only a little trickier than a normal network device configuration, you need to remember to bring up the FRAD device before the DLCI encapsulation devices.

     # Configure the frad hardware and the DLCI parameters
     /sbin/fradcfg /etc/frad/router.conf || exit 1
     /sbin/dlcicfg file /etc/frad/router.conf
     #
     # Bring up the FRAD device
     ifconfig sdla0 up
     #
     # Configure the DLCI encapsulation interfaces and routing
     ifconfig dlci00 192.168.10.1 pointopoint 192.168.10.2 up
     route add -net 192.168.10.0 netmask 255.255.255.0 dlci00
     #
     ifconfig dlci01 192.168.11.1 pointopoint 192.168.11.2 up
     route add -net 192.168.11.0 netmask 255.255.255.0 dlci00
     #
     route add default dev dlci00
     #

6.10 IP Accounting

The IP accounting features of the Linux kernel allow you to collect and analyse some network usage data. The data collected comprises the number of packets and the number of bytes accumulated since the figures were last reset. You may specify a variety of rules to categorise the figures to suit whatever purpose you may have.

Kernel Compile Options:

     Networking options  --->
         [*] IP: accounting

After you have compiled and installed the kernel you need to use the ipfwadm command to configure IP accounting. There are many different ways of breaking down the accounting information that you might choose. I’ve picked a simple example of what might be useful to use, you should read the ipfwadm man page for more information.

Scenario: You have a ethernet network that is linked to the internet via a PPP link. On the ethernet you have a machine that offers a number of services and that you are interested in knowing how much traffic is generated by each of telnet, rlogin, ftp and world wide web traffic.

You might use a command set that looks like the following:

#
# Flush the accounting rules
ipfwadm -A -f
#
# Add rules for local ethernet segment
ipfwadm -A in -a -P tcp -D 44.136.8.96/29 20
ipfwadm -A out -a -P tcp -S 44.136.8.96/29 20
ipfwadm -A in -a -P tcp -D 44.136.8.96/29 23
ipfwadm -A out -a -P tcp -S 44.136.8.96/29 23
ipfwadm -A in -a -P tcp -D 44.136.8.96/29 80
ipfwadm -A out -a -P tcp -S 44.136.8.96/29 80
ipfwadm -A in -a -P tcp -D 44.136.8.96/29 513
ipfwadm -A out -a -P tcp -S 44.136.8.96/29 513
ipfwadm -A in -a -P tcp -D 44.136.8.96/29
ipfwadm -A out -a -P tcp -D 44.136.8.96/29
ipfwadm -A in -a -P udp -D 44.136.8.96/29
ipfwadm -A out -a -P udp  -D 44.136.8.96/29
ipfwadm -A in -a -P icmp -D 44.136.8.96/29
ipfwadm -A out -a -P icmp -D 44.136.8.96/29
#
# Rules for default
ipfwadm -A in -a -P tcp -D 0/0 20
ipfwadm -A out -a -P tcp -S 0/0 20
ipfwadm -A in -a -P tcp -D 0/0 23
ipfwadm -A out -a -P tcp -S 0/0 23
ipfwadm -A in -a -P tcp -D 0/0 80
ipfwadm -A out -a -P tcp -S 0/0 80
ipfwadm -A in -a -P tcp -D 0/0 513
ipfwadm -A out -a -P tcp -S 0/0 513
ipfwadm -A in -a -P tcp -D 0/0
ipfwadm -A out -a -P tcp -D 0/0
ipfwadm -A in -a -P udp -D 0/0
ipfwadm -A out -a -P udp  -D 0/0
ipfwadm -A in -a -P icmp -D 0/0
ipfwadm -A out -a -P icmp -D 0/0
#
# List the rules
ipfwadm -A -l -n
#

The last command lists each of the Accounting rules and displays the collected totals.

An important point to note when analysing IP accounting is that totals for all rules that match will be incremented so that to obtain differential figures you need to perform appropriate maths. For example if I wanted to know how much data was not ftp, telnet, rlogin or www I would substract the individual totals from the rule that matches all ports.

# ipfwadm -A -l -n
IP accounting rules
pkts bytes dir prot source               destination          ports
0     0 in  tcp  0.0.0.0/0            44.136.8.96/29       * -> 20
0     0 out tcp  44.136.8.96/29       0.0.0.0/0            20 -> *
0     0 in  tcp  0.0.0.0/0            44.136.8.96/29       * -> 23
0     0 out tcp  44.136.8.96/29       0.0.0.0/0            23 -> *
10  1166 in  tcp  0.0.0.0/0            44.136.8.96/29       * -> 80
10   572 out tcp  44.136.8.96/29       0.0.0.0/0            80 -> *
242  9777 in  tcp  0.0.0.0/0            44.136.8.96/29       * -> 513
220 18198 out tcp  44.136.8.96/29       0.0.0.0/0            513 -> *
252 10943 in  tcp  0.0.0.0/0            44.136.8.96/29       * -> *
231 18831 out tcp  0.0.0.0/0            44.136.8.96/29       * -> *
0     0 in  udp  0.0.0.0/0            44.136.8.96/29       * -> *
0     0 out udp  0.0.0.0/0            44.136.8.96/29       * -> *
0     0 in  icmp 0.0.0.0/0            44.136.8.96/29       *
0     0 out icmp 0.0.0.0/0            44.136.8.96/29       *
0     0 in  tcp  0.0.0.0/0            0.0.0.0/0            * -> 20
0     0 out tcp  0.0.0.0/0            0.0.0.0/0            20 -> *
0     0 in  tcp  0.0.0.0/0            0.0.0.0/0            * -> 23
0     0 out tcp  0.0.0.0/0            0.0.0.0/0            23 -> *
10  1166 in  tcp  0.0.0.0/0            0.0.0.0/0            * -> 80
10   572 out tcp  0.0.0.0/0            0.0.0.0/0            80 -> *
243  9817 in  tcp  0.0.0.0/0            0.0.0.0/0            * -> 513
221 18259 out tcp  0.0.0.0/0            0.0.0.0/0            513 -> *
253 10983 in  tcp  0.0.0.0/0            0.0.0.0/0            * -> *
231 18831 out tcp  0.0.0.0/0            0.0.0.0/0            * -> *
0     0 in  udp  0.0.0.0/0            0.0.0.0/0            * -> *
0     0 out udp  0.0.0.0/0            0.0.0.0/0            * -> *
0     0 in  icmp 0.0.0.0/0            0.0.0.0/0            *
0     0 out icmp 0.0.0.0/0            0.0.0.0/0            *
#

6.11 IP Aliasing

There are some applications where being able to configure multiple IP addresses to a single network device is useful. Internet Service Providers often use this facility to provide a ‘customised’ to their World Wide Web and ftp offerings for their customers.

Kernel Compile Options:

     Networking options  --->
         ....
         [*] Network aliasing
         ....
         <*> IP: aliasing support

After compiling and installing your kernel with IP Alias support configuration is very simple. The aliases are added to virtual network devices associated with the actual network device. A simple naming convention applies to these devices being <devname>:<virtual dev num>, e.g. eth0:0, ppp0:10 etc.

For example, assume you have an ethernet network that supports two different IP subnetworks simultaneously and you wish your machine to have direct access to both, you could use something like:

     #
     # ifconfig eth0:0 192.168.1.1 netmask 255.255.255.0 up
     # route add -net 192.168.1.0 netmask 255.255.255.0 eth0:0
     #
     # ifconfig eth0:1 192.168.10.1 netmask 255.255.255.0 up
     # route add -net 192.168.10.0 netmask 255.255.255.0 eth0:0
     #

To delete an alias you simply add a ‘-’ to the end of its name and refer to it and is as simple as:

     # ifconfig eth0:0- 0

All routes associated with that alias will also be deleted automatically.

6.12 IP Firewall

IP Firewall and Firewalling issues are covered in more depth in the Firewall-HOWTO <Firewall-HOWTO.html>. IP Firewalling allows you to secure your machine against unauthorised network access by filtering or allowing datagrams from or to IP addresses that you nominate. There are three different classes of rules, incoming filtering, outgoing filtering and forwarding filtering. Incoming rules are applied to datagrams that are received by a network device. Outgoing rules are applied to datagrams that are to be transmitted by a network device. Forwarding rules are applied to datagrams that are received and are not for this machine, ie datagrams that would be routed.

Kernel Compile Options:

     Networking options  --->
         [*] Network firewalls
         ....
         [*] IP: forwarding/gatewaying
         ....
         [*] IP: firewalling
         [ ] IP: firewall packet logging

Configuration of the IP firewall rules is performed using the ipfwadm command. As I mentioned earlier, security is not something I am expert at, so while I will present an example you can use, you should do your own research and develop your own rules if security is important to you.

Probably the most common use of IP firewall is when you are using your linux machine as a router and firewall gateway to protect your local network from unauthorised access from outside your network.

The following configuration is based on a contribution from Arnt Gulbrandsen, <agulbra@troll.no>.

The example describes the configuration of the firewall rules on the Linux firewall/router machine illustrated in this diagram:

     -                                   -
      \                                  | 172.16.37.0
       \                                 |   /255.255.255.0
        \                 ---------      |
         |  172.16.174.30 | Linux |      |
     NET =================|  f/w  |------|    ..37.19
         |    PPP         | router|      |  --------
        /                 ---------      |--| Mail |
       /                                 |  | /DNS |
      /                                  |  --------
     -                                   -

The following commands would normally be placed in an rc file so that they were automatically started each time the system boots. For maximum security they would be performed after the network interfaces are configured, but before the interfaces are actually brought up to prevent anyone gaining access while the firewall machine is rebooting.

     #!/bin/sh
     # Flush the 'Forwarding' rules table
     # Change the default policy to 'accept'
     #
     /sbin/ipfwadm -F -f
     /sbin/ipfwadm -F -p accept
     #
     # .. and for 'Incoming'
     #
     /sbin/ipfwadm -I -f
     /sbin/ipfwadm -I -p accept
     # First off, seal off the PPP interface
     # I'd love to use '-a deny' instead of '-a reject -y' but then it
     # would be impossible to originate connections on that interface too.
     # The -o causes all rejected datagrams to be logged. This trades
     # disk space against knowledge of an attack of configuration error.
     #
     /sbin/ipfwadm -I -a reject -y -o -P tcp -S 0/0 -D 172.16.174.30

# Throw away certain kinds of obviously forged packets right away:
# Nothing should come from multicast/anycast/broadcast addresses
#
/sbin/ipfwadm -F -a deny -o -S 224.0/3 -D 172.16.37.0/24
#
# and nothing coming from the loopback network should ever be
# seen on a wire
#
/sbin/ipfwadm -F -a deny -o -S 127.0/8 -D 172.16.37.0/24
# accept incoming SMTP and DNS connections, but only
# to the Mail/Name Server
#
/sbin/ipfwadm -F -a accept -P tcp -S 0/0 -D 172.16.37.19 25 53
#
# DNS uses UDP as well as TCP, so allow that too
# for questions to our name server
#
/sbin/ipfwadm -F -a accept -P udp -S 0/0 -D 172.16.37.19 53
#
# but not "answers" coming to dangerous ports like NFS and
# Larry McVoy's NFS extension.  If you run squid, add its port here.
#
/sbin/ipfwadm -F -a deny -o -P udp -S 0/0 53 \
        -D 172.16.37.0/24 2049 2050
# answers to other user ports are okay
#
/sbin/ipfwadm -F -a accept -P udp -S 0/0 53 \
        -D 172.16.37.0/24 53 1024:65535
# Reject incoming connections to identd
# We use 'reject' here so that the connecting host is told
# straight away not to bother continuing, otherwise we'd experience
# delays while ident timed out.
#
/sbin/ipfwadm -F -a reject -o -P tcp -S 0/0 -D 172.16.37.0/24 113
# Accept some common service connections from the 192.168.64 and
# 192.168.65 networks, they are friends that we trust.
#
/sbin/ipfwadm -F -a accept -P tcp -S 192.168.64.0/23 \
        -D 172.16.37.0/24 20:23
# accept and pass through anything originating inside
#
/sbin/ipfwadm -F -a accept -P tcp -S 172.16.37.0/24 -D 0/0

     # deny most other incoming TCP connections and log them
     # (append 1:1023 if you have problems with ftp not working)
     #
     /sbin/ipfwadm -F -a deny -o -y -P tcp -S 0/0 -D 172.16.37.0/24
     # ... for UDP too
     #
     /sbin/ipfwadm -F -a deny -o -P udp -S 0/0 -D 172.16.37.0/24

Good firewall configurations are a little tricky. This example should be a reasonable starting point for you. The ipfwadm manual page offers some assistance in how to use the tool. If you intend to configure a firewall, be sure to ask around and get as much advice from sources you consider reliable and get someone to test/sanity check your configuration from the outside.

6.13 IPIP Encapsulation

Why would you want to encapsulate IP datagrams within IP datagrams? It must seem an odd thing to do if you’ve never seen an application of it before. Ok, here are a couple of common places where it is used: Mobile-IP and IP-Multicast. What is perhaps the most widely spread use of it though is also the least well known, Amateur Radio.

Kernel Compile Options:

     Networking options  --->
         [*] TCP/IP networking
         [*] IP: forwarding/gatewaying
         ....
         <*> IP: tunneling

IP tunnel devices are called ‘tunl0’, ‘tunl1’ etc.

”But why ?”. Ok, ok. Conventional IP routing rules mandate that an IP network comprises a network address and a network mask. This produces a series of contiguous addresses that may all be routed via a single routing entry. This is very convenient, but it means that you may only use any particular IP address while you are connected to the particular piece of network to which it belongs. In most instances this is ok, but if you are a mobile netizen then you may not be able to stay connected to the one place all the time. IP/IP encapsulation (IP tunneling) allows you to overcome this restriction by allowing datagrams destined for your IP address to be wrapped up and redirected to another IP address. If you know that you’re going to be operating from some other IP network for some time you can set up a machine on your home network to accept datagrams to your IP address and redirect them to the address that you will actually be using temporarily.

6.13.1 A tunneled network configuration.

As always, I believe a diagram will save me lots of confusing text, so here is one:

      192.168.1/24                          192.168.2/24
          -                                     -
          |      ppp0 =            ppp0 =       |
          |  aaa.bbb.ccc.ddd  fff.ggg.hhh.iii   |
          |                                     |
          |   /-----\                 /-----\   |
          |   |     |       //        |     |   |
          |---|  A  |------//---------|  B  |---|
          |   |     |     //          |     |   |
          |   \-----/                 \-----/   |
          |                                     |
          -                                     -

The diagram illustrates another possible reason to use IPIP encapsulation, virtual private networking. This example presupposes that you have two machines each with a simple dial up internet connection. Each host is allocated just a single IP address. Behind each of these machines are some private local area networks configured with reserved IP network addresses. Suppose that you want to allow any host on network A to connect to any host on network B, just as if they were properly connected to the Internet with a network route. IPIP encapsulation will allow you to do this. Note, encapsulation does not solve the problem of how you get the hosts on networks A and B to talk to any other on the Internet, you still need tricks like IP Masquerade for that. Encapsulation is normally performed by machine functioning as routers.

Linux router ‘A’ would be configured with:

     #
     PATH=/sbin:/usr/sbin
     #
     # Ethernet configuration
     ifconfig eth0 192.168.1.1 netmask 255.255.255.0 up
     route add -net 192.168.1.0 netmask 255.255.255.0 eth0
     #
     # ppp0 configuration (start ppp link, set default route)
     pppd
     route add default ppp0
     #
     # Tunnel device configuration
     ifconfig tunl0 192.168.1.1 up
     route add -net 192.168.2.0 netmask 255.255.255.0 gw fff.ggg.hhh.iii tunl0

Linux router ‘B’ would be configured with:

     #
     PATH=/sbin:/usr/sbin
     #
     # Ethernet configuration
     ifconfig eth0 192.168.2.1 netmask 255.255.255.0 up
     route add -net 192.168.2.0 netmask 255.255.255.0 eth0

     #
     # ppp0 configuration (start ppp link, set default route)
     pppd
     route add default ppp0
     #
     # Tunnel device configuration
     ifconfig tunl0 192.168.2.1 up
     route add -net 192.168.1.0 netmask 255.255.255.0 gw aaa.bbb.ccc.ddd tunl0

The command:

     route add -net 192.168.1.0 netmask 255.255.255.0 gw aaa.bbb.ccc.ddd tunl0

reads: ‘Send any datagrams destined for 192.168.1.0/24 inside an IPIP encap datagram with a destination address of aaa.bbb.ccc.ddd’.

Note that the configurations are reciprocated at either end. The tunnel device uses the ‘gw’ in the route as the destination of the IP datagram in which it will place the datagram it has received to route. That machine must know how to decapsulate IPIP datagrams, that is, it must also be configured with a tunnel device.

6.13.2 A tunneled host configuration.

It doesn’t have to be a whole network you route. You could for example route just a single IP address. In that instance you might configure the tunl device on the ‘remote’ machine with its home IP address and at the A end just use a host route (and Proxy Arp) rather than a network route via the tunnel device. Let’s redraw and modify our configuration appropriately. Now we have just host ‘B’ which to want to act and behave as if it is both fully connected to the Internet and also part of the remote network supported by host ‘A’:

      192.168.1/24
          -
          |      ppp0 =                ppp0 =
          |  aaa.bbb.ccc.ddd      fff.ggg.hhh.iii
          |
          |   /-----\                 /-----\
          |   |     |       //        |     |
          |---|  A  |------//---------|  B  |
          |   |     |     //          |     |
          |   \-----/                 \-----/
          |                      also: 192.168.1.12
          -

Linux router ‘A’ would be configured with:

     #

PATH=/sbin:/usr/sbin
#
# Ethernet configuration
ifconfig eth0 192.168.1.1 netmask 255.255.255.0 up
route add -net 192.168.1.0 netmask 255.255.255.0 eth0
#
# ppp0 configuration (start ppp link, set default route)
pppd
route add default ppp0
#
# Tunnel device configuration
ifconfig tunl0 192.168.1.1 up
route add -host 192.168.1.12 gw fff.ggg.hhh.iii tunl0
#
# Proxy ARP for the remote host
arp -s 192.168.1.12 xx:xx:xx:xx:xx:xx pub

Linux host ‘B’ would be configured with:

#
PATH=/sbin:/usr/sbin
#
# ppp0 configuration (start ppp link, set default route)
pppd
route add default ppp0
#
# Tunnel device configuration
ifconfig tunl0 192.168.1.12 up
route add -net 192.168.1.0 netmask 255.255.255.0 gw aaa.bbb.ccc.ddd tunl0

This sort of configuration is more typical of a Mobile-IP application. Where a single host wants to roam around the Internet and maintain a single usable IP address the whole time. You should refer to the Mobile-IP section for more information on how that is handled in practice.

6.14 IPX (AF IPX)

The IPX protocol is most commonly utilised in Novell NetWare(tm) local area network environments. Linux includes support for this protocol and may be configured to act as a network endpoint, or as a router for IPX.

Kernel Compile Options:

Networking options  --->
    [*] The IPX protocol
    [ ] Full internal IPX network

The IPX protocol and the NCPFS are covered in greater depth in the IPX-HOWTO <IPX-HOWTO.html>.

6.15 IPv6

Just when you thought you were beginning to understand IP networking the rules get changed! IPv6 is the shorthand notation for version 6 of the Internet Protocol. IPv6 was developed primarily to overcome the concerns in the Internet community that there would soon be a shortage of IP addresses to allocate. IPv6 addresses are 16 bytes long (128 bits). IPv6 incorporates a number of other changes, mostly simplifications, that will make IPv6 networks more managable than IPv4 networks.

Linux already has a working, but not completed, IPv6 implementation in the 2.1.* series kernels.

If you wish to experiment with this next generation Internet technology, or have a requirement for it, then you should read the IPv6-FAQ which is available from www.terra.net <http://www.terra.net/ipv6/>.

6.16 ISDN

The Integrated Services Digital Network (ISDN) is a series of standards that specify a general purpose switched digital data network. An ISDN ‘call’ creates a synchronous point to point data service to the destination. ISDN is generally delivered on a high speed link that is broken down into a number of discrete channels. There are two different types of channels, the ‘B Channels’ which will actually carry the user data and a single channel called the ‘D channel’ which is used to send control information to the ISDN exchange to establish calls and other functions. In Australia for example, ISDN may be delivered on a 2Mbps link that is broken into 30 discreet 64kbps B channels with one 64kbps D channel. Any number of channels may be used at a time and in any combination. You could for example establish 30 seperate calls to 30 different destinations at 64kbps each, or you could establish 15 calls to 15 different destinations at 128kbps each (two channels used per call), or just a small number of calls and leave the rest idle. A channel may be used for either incoming or outgoing calls. The original intention of ISDN was to allow Telecommunications companies to provide a single data service which could deliver either telephone (via digitised voice) or data services to your home or business without requiring you to make any special configuration changes.

There are a few different ways to connect your computer to an ISDN service. One way is to use a device called a ‘Terminal Adaptor’ which plugs into the Network Terminating Unit that you telecommunications carrier will have installed when you got your ISDN service and presents a number of serial interfaces. One of those interfaces is used to enter commands to establish calls and configuration and the others are actually connected to the network devices that will use the data circuits when they are established. Linux will work in this sort of configuration without modification, you just treat the port on the Terminal Adaptor like you would treat any other serial device. Another way, which is the way the kernel ISDN support is designed for allows you to install an ISDN card into your Linux machine and then has your Linux software handle the protocols and make the calls itself.

Kernel Compile Options:

     ISDN subsystem  --->
             <*> ISDN support
             [ ] Support synchronous PPP
             [ ] Support audio via ISDN
             < > ICN 2B and 4B support
             < > PCBIT-D support

             < > Teles/NICCY1016PC/Creatix support

The Linux implementation of ISDN support a number of different types of internal ISDN cards. These are those listed in the kernel configuration options:

• ICN 2B and 4B

• Octal PCBIT-D

• Teles ISDN-cards and compatibles

Some of these cards require software to be downloaded to them to make them operational. There is a seperate utility to do this with.

Full details on how to configure the Linux ISDN support is available from the /usr/src/linux/Documentation/isdn/ directory and an FAQ dedicated to isdn4linux is available at

www.lrz-muenchen.de <http://www.lrz-muenchen.de/ ui161ab/www/isdn/>. (You can click on the

english flag to get an english version).

A note about PPP. The PPP suite of protocols will operate over either asynchronous or synchronous serial lines. The commonly distributed PPP daemon for Linux ‘pppd’ supports only asynchronous mode. If you wish to run the PPP protocols over your ISDN service you need a specially modified version. Details of where to find it are available in the documentation referred to above.

6.17 IP Masquerade

Many people have a simple dialup account to connect to the Internet. Nearly everybody using this sort of configuration is allocated a single IP address by the Internet Service Provider. This is normally enough to allow only one host full access to the network. IP Masquerade is a clever trick that enables you to have many machines make use of that one IP address, by causing the other hosts to look like, hence the term masquerade, the machine supporting the dialup connection. There is a small caveat and that is that the masqerade function nearly always works only in one direction, that is the masqueraded hosts can make calls out, but they cannot accept or receive network connections from remote hosts. This means that some network services do not work such as talk and others such as ftp must be configured to operate in passive (PASV) mode to operate. Fortunately the most common network services such as telnet, World Wide Web and irc do work just fine.

Kernel Compile Options:

     Code maturity level options  --->
         [*] Prompt for development and/or incomplete code/drivers
     Networking options  --->
         [*] Network firewalls
         ....
         [*] TCP/IP networking
         [*] IP: forwarding/gatewaying
         ....
         [*] IP: masquerading (EXPERIMENTAL)

Normally you have your linux machine supporting a slip or PPP dialup line just as it would if it were a standalone machine. Additionally it would have another network device configured, perhaps an ethernet, configured with one of the reserved network addresses. The hosts to be masqueraded would be on this second network. Each of these hosts would have the IP address of the ethernet port of the linux machine set as their default gateway or router.

A typical configuration might look something like this:

     -                                   -
      \                                  | 192.168.1.0
       \                                 |   /255.255.255.0
        \                 ---------      |
         |                | Linux | .1.1 |
     NET =================| masq  |------|
         |    PPP/slip    | router|      |  --------
        /                 ---------      |--| host |
       /                                 |  |      |
      /                                  |  --------
     -                                   -

The most relevant commands for this configuration are:

     # Network route for ethernet
     route add -net 192.168.1.0 netmask 255.255.255.0 eth0
     #
     # Default route to the rest of the internet.
     route add default ppp0
     #
     # Cause all hosts on the 192.168.1/24 network to be masqueraded.
     ipfwadm -F -a m -S 192.168.1.0/24 -D 0.0.0.0/0

You can get more information on the Linux IP Masquerade feature from the IP Masquerade Resource Page

<http://www.hwy401.com/achau/ipmasq/>

6.18 IP Transparent Proxy

IP transparent proxy is a feature that enables you to redirect servers or services destined for another machine to those services on this machine. Typically this would be useful where you have a linux machine as a router and also provides a proxy server. You would redirect all connections destined for that service remotely to the local proxy server.

Kernel Compile Options:

     Code maturity level options  --->
             [*] Prompt for development and/or incomplete code/drivers
     Networking options  --->
             [*] Network firewalls
             ....

             [*] TCP/IP networking
             ....
             [*] IP: firewalling
             ....
             [*] IP: transparent proxy support (EXPERIMENTAL)

Configuration of the transparent proxy feature is performed using the ipfwadm command.

An example that might be useful is as follows:

     ipfwadm -I -a accept -D 0/0 80 -r 8080

This example will cause any connection attempts to port 80 (www) on any host to be redirected to port 8080 on this host. This could be used to ensure that all WWW traffic from your network is automatically directed to a local WWW cache program.

6.19 Mobile IP

The term ”IP mobility” describes the ability of a host that is able to move its network connection from one point on the Internet to another without changing its IP address or losing connectivity. Usually when an IP host changes its point of connectivity it must also change its IP address. IP Mobility overcomes this problem by allocating a fixed IP address to the mobile host and using IP encapsulation (tunneling) with automatic routing to ensure that datagrams destined for it are routed to the actual IP address it is currently using.

A project is underway to provide a complete set of IP mobility tools for Linux. The Status of the project and tools may be obtained from the: Linux Mobile IP Home Page

<http://anchor.cs.binghamton.edu/ mobileip/>.

6.20 Multicast

IP Multicast allows an arbitrary number of IP hosts on disparate IP networks to have IP datagrams simultaneously routed to them. This mechanism is exploited to provide Internet wide ”broadcast” material such as audio and video transmissions and other novel applications.

Kernel Compile Options:

     Networking options  --->
             [*] TCP/IP networking
             ....
             [*] IP: multicasting

A suite of tools and some minor network configuration is required. One source of information on how to install and configure these for Linux is provided at www.teksouth.com

<http://www.teksouth.com/linux/multicast/>.

6.21 NAT - Network Address Translation

The IP Network Address Translation facility is pretty much the standardised big brother of the Linux IP Masquerade facility. It is specified in some detail in RFC-1631 at your nearest RFC archive. NAT provides features that IP-Masquerade does not that make it eminently more suitable for use in corporate firewall router designs and larger scale installations.

An alpha implementation of NAT for Linux 2.0.29 kernel has been developed by Michael.Hasenstein,

Michael.Hasenstein@informatik.tu-chemnitz.de. Michaels documentation and implementation are

available from:

Linux IP Network Address Web Page

<http://www.csn.tu-chemnitz.de/HyperNews/get/linux-ip-nat.html>

Newer Linux 2.1.* kernels also include some NAT functionality in the routing algorithm.

6.22 NetRom (AF NETROM)

NetRom device names are ‘nr0’, ‘nr1’, etc.

Kernel Compile Options:

     Networking options  --->
         [*] Amateur Radio AX.25 Level 2
         [*] Amateur Radio NET/ROM

The AX25, Netrom and Rose protocols are covered by the AX25-HOWTO <AX25-HOWTO.html>. These protocols are used by Amateur Radio Operators world wide in packet radio experimentation.

Most of the work for implementation of these protocols has been done by Jonathon Naylor,

jsn@cs.nott.ac.uk.

6.23 PLIP

PLIP device names are ‘plip0’, ‘plip1 and plip2.

Kernel Compile Options:

     Networking options  --->
         <*> PLIP (parallel port) support

plip (Parallel Line IP), is like SLIP, in that it is used for providing a point to point network connection between two machines, except that it is designed to use the parallel printer ports on your machine instead of the serial ports (a cabling diagram in included in the cabling diagram section later in this document). Because it is possible to transfer more than one bit at a time with a parallel port, it is possible to attain higher speeds with the plip interface than with a standard serial device. In addition, even the simplest of parallel ports, printer ports, can be used in lieu of you having to purchase comparatively expensive 16550AFN UART’s for your serial ports. PLIP uses a lot of CPU compared to a serial link and is most certainly not a good option if you can obtain some cheap ethernet cards, but it will work when nothing else is available and will work quite well. You should expect a data transfer rate of about 20 kilobytes per second when a link is running well.

The PLIP device drivers competes with the parallel device driver for the parallel port hardware. If you wish to use both drivers then you should compile them both as modules to ensure that you are able to select which port you want to use for PLIP and which ports you want for the printer driver. Refer to the Modules-HOWTO <Modules-HOWTO.html> for more information on kernel module configuration.

Please note that some laptops use chipsets that will not work with PLIP because they do not allow some combinations of signals that PLIP relies on, that printers don’t use.

The Linux plip interface is compatible with the Crynwyr Packet Driver PLIP and this will mean that you can connect your Linux machine to a DOS machine running any other sort of tcp/ip software via plip.

In the 2.0.* series kernel the plip devices are mapped to i/o port and IRQ as follows:

     device  i/o     IRQ
     ------  -----   ---
     plip0   0x3bc   5
     plip1   0x378   7
     plip2   0x278   2

If your parallel ports don’t match any of the above combinations then you can change the IRQ of a port using the ifconfig command using the ‘irq’ parameter. Be sure to enable IRQ’s on your printer ports in your ROM BIOS if it supports this option.

In late 2.1.* series kernel with Plug’n’Play support the plip devices are allocated sequentially as they are detected just like the ethernet devices with plip0 being the first allocated.

When compiling the kernel, there is only one file that might need to be looked at to configure plip. That file is /usr/src/linux/driver/net/CONFIG and it contains plip timers in milliseconds. The defaults are probably ok in most cases. You will probably need to increase them if you have an especially slow computer, in which case the timers to increase are actually on the other computer. A program called plipconfig exists that allows you to change these timer settings without recompiling your kernel. It is supplied with many Linux distributions.

To configure a plip interface, you will need to add the following lines to your network rc file:

     #
     # Attach a PLIP interface
     #
     #  configure first parallel port as a plip device
     /sbin/ifconfig plip0 IPA.IPA.IPA.IPA pointopoint IPR.IPR.IPR.IPR up
     #
     # End plip

Where:

IPA.IPA.IPA.IPA

represents your IP address.

IPR.IPR.IPR.IPR represents the IP address of the remote machine.

The pointopoint parameter has the same meaning as for SLIP, in that it specifies the address of the machine at the other end of the link.

In almost all respects you can treat a plip interface as though it were a SLIP interface, except that neither dip nor slattach need be, nor can be, used.

Further information on PLIP may be obtained from the PLIP-mini-HOWTO <mini/PLIP>

6.24 PPP

PPP devices names are ‘ppp0’, ‘ppp1, etc. Devices are numbered sequentially with the first device configured receiving ‘0’.

Kernel Compile Options:

     Networking options  --->
         <*> PPP (point-to-point) support

PPP configuration is covered in detail in the PPP-HOWTO <PPP-HOWTO.html>.

6.24.1 Maintaining a permanent connection to the net with pppd.

If you are fortunate enough to have a semi permanent connection to the net and would like to have your machine automatically redial your PPP connection if it is lost then here is a simple trick to do so.

Configure PPP such that it can be started by the root user by issuing the command:

     # pppd

Be sure that you have the ‘-detach’ option configured in your /etc/ppp/options file. Then, insert the following line into your /etc/inittab file, down with the getty definitions:

     pd:23:respawn:/usr/sbin/pppd

This will cause the init program to spawn and monitor the pppd program and automatically restart it if it dies.

6.25 Rose protocol (AF ROSE)

Rose device names are ‘rs0’, ‘rs1’, etc. in 2.1.* kernels. Rose is available in the 2.1.* kernels.

Kernel Compile Options:

     Networking options  --->
         [*] Amateur Radio AX.25 Level 2
         <*> Amateur Radio X.25 PLP (Rose)

The AX25, Netrom and Rose protocols are covered by the AX25-HOWTO <AX25-HOWTO.html>. These protocols are used by Amateur Radio Operators world wide in packet radio experimentation.

Most of the work for implementation of these protocols has been done by Jonathon Naylor,

jsn@cs.nott.ac.uk.

6.26 SAMBA - ‘NetBEUI’, ‘NetBios’ support.

SAMBA is an implementation of the Session Management Block protocol. Samba allows Microsoft and other systems to mount and use your disks and printers.

SAMBA and its configuration are covered in detail in the SMB-HOWTO <SMB-HOWTO.html>.

6.27 SLIP client

SLIP devices are named ‘sl0’, ‘sl1’ etc. with the first device configured being assigned ‘0’ and the rest incrementing sequentially as they are configured.

Kernel Compile Options:

     Network device support  --->
         [*] Network device support
         <*> SLIP (serial line) support
         [ ]  CSLIP compressed headers
         [ ]  Keepalive and linefill
         [ ]  Six bit SLIP encapsulation

SLIP (Serial Line Internet Protocol) allows you to use tcp/ip over a serial line, be that a phone line with a dialup modem, or a leased line of some sort. Of course to use SLIP you need access to a SLIP-server in your area. Many universities and businesses provide SLIP access all over the world.

Slip uses the serial ports on your machine to carry IP datagrams. To do this it must take control of the serial device. Slip device names are named sl0 , sl1 etc. How do these correspond to your serial devices ? The networking code uses what is called an ioctl (i/o control) call to change the serial devices into SLIP devices. There are two programs supplied that can do this, they are called dip and slattach

6.27.1 dip

dip (Dialup IP) is a smart program that is able to set the speed of the serial device, command your modem to dial the remote end of the link, automatically log you into the remote server, search for messages sent to you by the server and extract information for them such as your IP address and perform the ioctl necessary to switch your serial port into SLIP mode. dip has a powerful scripting ability and it is this that you can exploit to automate your logon procedure.

You can find it at: sunsite.unc.edu

<ftp://sunsite.unc.edu/pub/Linux/system/Network/serial/dip/dip337o-uri.tgz>.

To install it, try the following:

     #
     # cd /usr/src
     # gzip -dc dip337o-uri.tgz | tar xvf -
     # cd dip-3.3.7o
     <edit Makefile>
     # make install
     #

The Makefile assumes the existence of a group called uucp, but you might like to change this to either dip or SLIP depending on your configuration.

6.27.2 slattach

slattach as contrasted with dip is a very simple program, that is very easy to use, but does not have the sophistication of dip. It does not have the scripting ability, all it does is configure your serial device as a SLIP device. It assumes you have all the information you need and the serial line is established before you invoke it. slattach is ideal to use where you have a permanent connection to your server, such as a physical cable, or a leased line.

6.27.3 When do I use which ?

You would use dip when your link to the machine that is your SLIP server is a dialup modem, or some other temporary link. You would use slattach when you have a leased line, perhaps a cable, between your machine and the server and there is no special action needed to get the link working. See section ‘Permanent Slip connection’ for more information.

Configuring SLIP is much like configuring an Ethernet interface (read section ‘Configuring an ethernet device’ above). However there are a few key differences.

First of all, SLIP links are unlike ethernet networks in that there is only ever two hosts on the network, one at each end of the link. Unlike an ethernet that is available for use as soon are you are cabled, with SLIP, depending on the type of link you have, you may have to initialize your network connection in some special way.

If you are using dip then this would not normally be done at boot time, but at some time later, when you were ready to use the link. It is possible to automate this procedure. If you are using slattach then you will probably want to add a section to your rc.inet1 file. This will be described soon.

There are two major types of SLIP servers: Dynamic IP address servers and static IP address servers. Almost every SLIP server will prompt you to login using a username and password when dialing in. dip can handle logging you in automatically.

6.27.4 Static SLIP server with a dialup line and DIP.

A static SLIP server is one in which you have been supplied an IP address that is exclusively yours. Each time you connect to the server, you will configure your SLIP port with that address. The static SLIP server will answer your modem call, possibly prompt you for a username and password, and then route any datagrams destined for your address to you via that connection. If you have a static server, then you may want to put entries for your hostname and IP address (since you know what it will be) into your /etc/hosts. You should also configure some other files such as: rc.inet2, host.conf, resolv.conf, /etc/HOSTNAME and rc.local. Remember that when configuring rc.inet1, you don’t need to add any special commands for your SLIP connection since it is dip that does all of the hard work for you in configuring your interface. You will need to give dip the appropriate information and it will configure the interface for you after commanding the modem to establish the call and logging you into your SLIP server.

If this is how your SLIP server works then you can move to section ‘Using Dip’ to learn how to configure dip appropriately.

6.27.5 Dynamic SLIP server with a dialup line and DIP.

A dynamic SLIP server is one which allocates you an IP address randomly, from a pool of addresses, each time you logon. This means that there is no guarantee that you will have any particular address each time, and that address may well be used by someone else after you have logged off. The network administrator who configured the SLIP server will have assigned a pool of address for the SLIP server to use, when the server receives a new incoming call, it finds the first unused address, guides the caller through the login process and then prints a welcome message that contains the IP address it has allocated and will proceed to use that IP address for the duration of that call.

Configuring for this type of server is similar to configuring for a static server, except that you must add a step where you obtain the IP address that the server has allocated for you and configure your SLIP device with that.

Again, dip does the hard work and new versions are smart enough to not only log you in, but to also be able to automatically read the IP address printed in the welcome message and store it so that you can have it configure your SLIP device with it.

If this is how your SLIP server works then you can move to section ‘Using Dip’ to learn how to configure dip appropriately.

6.27.6 Using DIP.

As explained earlier, dip is a powerful program that can simplify and automate the process of dialing into the SLIP server, logging you in, starting the connection and configuring your SLIP devices with the appropriate ifconfig and route commands.

Essentially to use dip you’ll write a ‘dip script’, which is basically a list of commands that dip understands that tell dip how to perform each of the actions you want it to perform. See sample.dip that comes supplied with dip to get an idea of how it works. dip is quite a powerful program, with many options.

Instead of going into all of them here you should look at the man page, README and sample files that will have come with your version of dip.

You may notice that the sample.dip script assumes that you’re using a static SLIP server, so you know what your IP address is beforehand. For dynamic SLIP servers, the newer versions of dip include a command you can use to automatically read and configure your SLIP device with the IP address that the dynamic server allocates for you. The following sample is a modified version of the sample.dip that came supplied with dip337j-uri.tgz and is probably a good starting point for you. You might like to save it as /etc/dipscript and edit it to suit your configuration:

#
# sample.dip    Dialup IP connection support program.
#
#               This file (should show) shows how to use the DIP
#       This file should work for Annex type dynamic servers, if you
#       use a static address server then use the sample.dip file that
#       comes as part of the dip337-uri.tgz package.
#
#
# Version:      @(#)sample.dip  1.40    07/20/93
#
# Author:       Fred N. van Kempen, <waltje@uWalt.NL.Mugnet.ORG>
#
main:
# Next, set up the other side's name and address.
# My dialin machine is called 'xs4all.hacktic.nl' (== 193.78.33.42)
get $remote xs4all.hacktic.nl
# Set netmask on sl0 to 255.255.255.0
netmask 255.255.255.0
# Set the desired serial port and speed.
port cua02
speed 38400
# Reset the modem and terminal line.
# This seems to cause trouble for some people!
reset
# Note! "Standard" pre-defined "errlevel" values:
#  0 - OK
#  1 - CONNECT
#  2 - ERROR
#
# You can change those grep'ping for "addchat()" in *.c...
# Prepare for dialing.
send ATQ0V1E1X4\r
wait OK 2

if $errlvl != 0 goto modem_trouble
dial 555-1234567
if $errlvl != 1 goto modem_trouble
# We are connected.  Login to the system.
login:
sleep 2
wait ogin: 20
if $errlvl != 0 goto login_trouble
send MYLOGIN\n
wait ord: 20
if $errlvl != 0 goto password_error
send MYPASSWD\n
loggedin:
# We are now logged in.
wait SOMEPROMPT 30
if $errlvl != 0 goto prompt_error
# Command the server into SLIP mode
send SLIP\n
wait SLIP 30
if $errlvl != 0 goto prompt_error
# Get and Set your IP address from the server.
#   Here we assume that after commanding the SLIP server into SLIP
#   mode that it prints your IP address
get $locip remote 30
if $errlvl != 0 goto prompt_error
# Set up the SLIP operating parameters.
get $mtu 296
# Ensure "route add -net default xs4all.hacktic.nl" will be done
default
# Say hello and fire up!
done:
print CONNECTED $locip ---> $rmtip
mode CSLIP
goto exit
prompt_error:
print TIME-OUT waiting for sliplogin to fire up...
goto error
login_trouble:
print Trouble waiting for the Login: prompt...
goto error

     password:error:
     print Trouble waiting for the Password: prompt...
     goto error
     modem_trouble:
     print Trouble occurred with the modem...
     error:
     print CONNECT FAILED to $remote
     quit
     exit:
     exit

The above example assumes you are calling a dynamic SLIP server, if you are calling a static SLIP server, then the sample.dip file that comes with dip337j-uri.tgz should work for you.

When dip is given the get $local command it searches the incoming text from the remote end for a string that looks like an IP address, ie strings numbers separated by ‘.’ characters. This modification was put in place specifically for dynamic SLIP servers, so that the process of reading the IP address granted by the server could be automated.

The example above will automatically create a default route via your SLIP link, if this is not what you want, you might have an ethernet connection that should be your default route, then remove the default command from the script. After this script has finished running, if you do an ifconfig command, you will see that you have a device sl0 . This is your SLIP device. Should you need to, you can modify its configuration manually, after the dip command has finished, using the ifconfig and route commands.

Please note that dip allows you to select a number of different protocols to use with the mode command, the most common example is cSLIP for SLIP with compression. Please note that both ends of the link must agree, so you should ensure that whatever you select agrees with what your server is set to.

The above example is fairly robust and should cope with most errors. Please refer to the dip man page for more information. Naturally you could, for example, code the script to do such things as redial the server if it doesn’t get a connection within a prescribed period of time, or even try a series of servers if you have access to more than one.

6.27.7 Permanent SLIP connection using a leased line and slattach.

If you have a cable between two machines, or are fortunate enough to have a leased line, or some other permanent serial connection between your machine and another, then you don’t need to go to all the trouble of using dip to set up your serial link. slattach is a very simple to use utility that will allow you just enough functionality to configure your connection.

Since your connection will be a permanent one, you will want to add some commands to your rc.inet1 file. In essence all you need to do for a permanent connection is ensure that you configure the serial device to the correct speed and switch the serial device into SLIP mode. slattach allows you to do this with one command. Add the following to your rc.inet1 file:

     #
     # Attach a leased line static SLIP connection
     #
     #  configure /dev/cua0 for 19.2kbps and cslip
     /sbin/slattach -p cslip -s 19200 /dev/cua0 &
     /sbin/ifconfig sl0 IPA.IPA.IPA.IPA pointopoint IPR.IPR.IPR.IPR up
     #
     # End static SLIP.

Where:

IPA.IPA.IPA.IPA represents your IP address.

IPR.IPR.IPR.IPR represents the IP address of the remote end.

slattach allocates the first unallocated SLIP device to the serial device specified. slattach starts with sl0 . Therefore the first slattach command attaches SLIP device sl0 to the serial device specified and sl1 the next time, etc.

slattach allows you to configure a number of different protocols with the -p argument. In your case you will use either SLIP or cSLIP depending on whether you want to use compression or not. Note: both ends must agree on whether you want compression or not.

6.28 SLIP server.

If you have a machine that is perhaps network connected, that you’d like other people be able to dial into and provide network services, then you will need to configure your machine as a server. If you want to use SLIP as the serial line protocol, then currently you have three options as to how to configure your Linux machine as a SLIP server. My preference would be to use the first presented, sliplogin, as it seems the easiest to configure and understand, but I will present a summary of each, so you can make your own decision.

6.28.1 Slip Server using sliplogin.

sliplogin is a program that you can use in place of the normal login shell for SLIP users that converts the terminal line into a SLIP line. It allows you to configure your Linux machine as either a static address server, users get the same address everytime they call in, or a dynamic address server, where users get an address allocated for them which will not necessarily be the same as the last time they called.

The caller will login as per the standard login process, entering their username and password, but instead of being presented with a shell after their login, sliplogin is executed which searches its configuration file (/etc/slip.hosts) for an entry with a login name that matches that of the caller. If it locates one, it configures the line as an 8bit clean line, and uses an ioctl call to convert the line discipline to SLIP. When this process is complete, the last stage of configuration takes place, where sliplogin invokes a shell script which configures the SLIP interface with the relevant ip address, netmask and sets appropriate routing in place. This script is usually called /etc/slip.login, but in a similar manner to getty, if you have certain callers that require special initialisation, then you can create configuration scripts called /etc/slip.login.loginname that will be run instead of the default specifically for them.

There are either three or four files that you need to configure to get sliplogin working for you. I will detail how and where to get the software and how each is configured in detail. The files are:

• /etc/passwd, for the dialin user accounts.

• /etc/slip.hosts, to contain the information unique to each dial-in user.

• /etc/slip.login, which manages the configuration of the routing that needs to be performed for the user.

• /etc/slip.tty, which is required only if you are configuring your server for dynamic address allocation and contains a table of addresses to allocate

• /etc/slip.logout, which contains commands to clean up after the user has hung up or logged out.

Where to get sliplogin You may already have the sliplogin package installed as part of your distribution, if not then sliplogin can be obtained from: sunsite.unc.edu

<ftp://sunsite.unc.edu/pub/linux/system/Network/serial/sliplogin-2.1.1.tar.gz>. The tar file

contains both source, precompiled binaries and a man page.

To ensure that only authorised users will be able to run sliplogin program, you should add an entry to your /etc/group file similar to the following:

      ..
     slip::13:radio,fred
      ..

When you install the sliplogin package, the Makefile will change the group ownership of the sliplogin program to slip, and this will mean that only users who belong to that group will be able to execute it. The example above will allow only users radio and fred to execute sliplogin.

To install the binaries into your /sbin directory and the man page into section 8, do the following:

     # cd /usr/src
     # gzip -dc .../sliplogin-2.1.1.tar.gz | tar xvf -
     # cd sliplogin-2.1.1
     # <..edit the Makefile if you don't use shadow passwords..>
     # make install

If you want to recompile the binaries before installation, add a make clean before the make install. If you want to install the binaries somewhere else, you will need to edit the Makefile install rule.

Please read the README files that come with the package for more information.

Configuring /etc/passwd for Slip hosts. Normally you would create some special logins for Slip callers in your /etc/passwd file. A convention commonly followed is to use the hostname of the calling host with a capital ‘S’ prefixing it. So, for example, if the calling host is called radio then you could create a /etc/passwd entry that looked like:

     Sradio:FvKurok73:1427:1:radio SLIP login:/tmp:/sbin/sliplogin

It doesn’t really matter what the account is called, so long as it is meaningful to you.

Note: the caller doesn’t need any special home directory, as they will not be presented with a shell from this machine, so /tmp is a good choice. Also note that sliplogin is used in place of the normal login shell.

Configuring /etc/slip.hosts The /etc/slip.hosts file is the file that sliplogin searches for entries matching the login name to obtain configuration details for this caller. It is this file where you specify the ip address and netmask that will be assigned to the caller and configured for their use. Sample entries for two hosts, one a static configuration for host radio and another, a dynamic configuration for user host albert might look like:

     #
     Sradio   44.136.8.99   44.136.8.100  255.255.255.0  normal      -1
     Salbert  44.136.8.99   DYNAMIC       255.255.255.0  compressed  60
     #

The /etc/slip.hosts file entries are:

1. the login name of the caller.

2. ip address of the server machine, ie this machine.

3. ip address that the caller will be assigned. If this field is coded DYNAMIC then an ip address will be allocated based on the information contained in your /etc/slip.tty file discussed later. Note: you must be using at least version 1.3 of sliplogin for this to work.

4. the netmask assigned to the calling machine in dotted decimal notation eg 255.255.255.0 for a Class C network mask.

5. the slip mode setting which allows you to enable/disable compression and slip other features. Allowable values here are ”normal” or ”compressed”.

6. a timeout parameter which specifies how long the line can remain idle (no datagrams received) before the line is automatically disconnected. A negative value disables this feature.

7. optional arguments.

Note: You can use either hostnames or IP addresses in dotted decimal notation for fields 2 and 3. If you use hostnames then those hosts must be resolvable, that is, your machine must be able to locate an ip address for those hostnames, otherwise the script will fail when it is called. You can test this by trying trying to telnet to the hostname, if you get the ‘Trying nnn.nnn.nnn...’ message then your machine has been able to find an ip address for that name. If you get the message ‘Unknown host’, then it has not. If not, either use ip addresses in dotted decimal notation, or fix up your name resolver configuration (See

section Name Resolution).

The most common slip modes are:

normal

to enable normal uncompressed SLIP.

compressed

to enable van Jacobsen header compression (cSLIP)

Naturally these are mutually exclusive, you can use one or the other. For more information on the other options available, refer to the man pages.

Configuring the /etc/slip.login file. After sliplogin has searched the /etc/slip.hosts and found a matching entry, it will attempt to execute the /etc/slip.login file to actually configure the SLIP interface with its ip address and netmask.

The sample /etc/slip.login file supplied with the sliplogin package looks like this:

     #!/bin/sh -
     #
     #       @(#)slip.login  5.1 (Berkeley) 7/1/90
     #
     # generic login file for a SLIP line.  sliplogin invokes this with
     # the parameters:
     #     $1       $2       $3    $4, $5, $6 ...
     #   SLIPunit ttyspeed   pid   the arguments from the slip.host entry
     #
     /sbin/ifconfig $1 $5 pointopoint $6 mtu 1500 -trailers up
     /sbin/route add $6
     arp -s $6 <hw_addr> pub
     exit 0
     #

You will note that this script simply uses the ifconfig and route commands to configure the SLIP device with its ipaddress, remote ip address and netmask and creates a route for the remote address via the SLIP device. Just the same as you would if you were using the slattach command.

Note also the use of Proxy ARP to ensure that other hosts on the same ethernet as the server machine will know how to reach the dial-in host. The <hw addr> field should be the hardware address of the ethernet card in the machine. If your server machine isn’t on an ethernet network then you can leave this line out completely.

Configuring the /etc/slip.logout file. When the call drops out, you want to ensure that the serial device is restored to its normal state so that future callers will be able to login correctly. This is achieved with the use of the /etc/slip.logout file. It is quite simple in format and is called with the same argument as the /etc/slip.login file.

     #!/bin/sh -
     #
     #               slip.logout
     #
     /sbin/ifconfig $1 down
     arp -d $6
     exit 0
     #

All it does is ‘down’ the interface which will delete the manual route previously created. It also uses the arp command to delete any proxy arp put in place, again, you don’t need the arp command in the script if your server machine does not have an ethernet port.

Configuring the /etc/slip.tty file. If you are using dynamic ip address allocation (have any hosts configured with the DYNAMIC keyword in the /etc/slip.hosts file, then you must configure the /etc/slip.tty file to list what addresses are assigned to what port. You only need this file if you wish your server to dynamically allocate addresses to users.

The file is a table that lists the tty devices that will support dial-in SLIP connections and the ip address that should be assigned to users who call in on that port.

Its format is as follows:

     # slip.tty    tty -> IP address mappings for dynamic SLIP
     # format: /dev/tty?? xxx.xxx.xxx.xxx
     #
     /dev/ttyS0      192.168.0.100
     /dev/ttyS1      192.168.0.101
     #

What this table says is that callers that dial in on port /dev/ttyS0 who have their remote address field in the /etc/slip.hosts file set to DYNAMIC will be assigned an address of 192.168.0.100.

In this way you need only allocate one address per port for all users who do not require an dedicated address for themselves. This helps you keep the number of addresses you need down to a minimum to avoid wastage.

6.28.2 Slip Server using dip.

Let me start by saying that some of the information below came from the dip man pages, where how to run Linux as a SLIP server is briefly documented. Please also beware that the following has been based on the dip337o-uri.tgz package and probably will not apply to other versions of dip. dip has an input mode of operation, where it automatically locates an entry for the user who invoked it and configures the serial line as a SLIP link according to information it finds in the /etc/diphosts file. This input mode of operation is activated by invoking dip as diplogin. This therefore is how you use dip as a SLIP server, by creating special accounts where diplogin is used as the login shell.

The first thing you will need to do is to make a symbolic link as follows:

     # ln -sf /usr/sbin/dip /usr/sbin/diplogin

You then need to add entries to both your /etc/passwd and your /etc/diphosts files. The entries you need to make are formatted as follows:

To configure Linux as a SLIP server with dip, you need to create some special SLIP accounts for users, where dip (in input mode) is used as the login shell. A suggested convention is that of having all SLIP accounts begin with a capital ‘S’, eg ‘Sfredm’.

A sample /etc/passwd entry for a SLIP user looks like:

     Sfredm:ij/SMxiTlGVCo:1004:10:Fred:/tmp:/usr/sbin/diplogin
     ^^^^^^^^^^^^^^
     |          |         |   |    |    |    \__diplogin as login shell
     |          |         |   |    |    \_______Home directory
     |          |         |   |    \____________User Full Name
     |          |         |   \_________________User Group ID
     |          |         \_____________________User ID
     |          \_______________________________Encrypted User Password
     \__________________________________________Slip User Login Name

After the user logs in, the login program, if it finds and verifies the user ok, will execute the diplogin command. dip, when invoked as diplogin knows that it should automatically assume that it is being used a login shell. When it is started as diplogin the first thing it does is use the getuid() function call to get the userid of whoever has invoked it. It then searches the /etc/diphosts file for the first entry that matches either the userid or the name of the tty device that the call has come in on and configures itself appropriately. By judicious decision as to whether to give a user an entry in the diphosts file, or whether to let the user be given the default configuration you can build your server in such a way that you can have a mix of static and dynamically assigned address users.

dip will automatically add a ‘Proxy-ARP’ entry if invoked in input mode, so you do not need to worry about manually adding such entries.

Configuring /etc/diphosts /etc/diphosts is used by dip to lookup preset configurations for remote hosts. These remote hosts might be users dialing into your linux machine, or they might be for machines that you dial into with your linux machine.

The general format for /etc/diphosts is as follows:

      ..
     Suwalt::145.71.34.1:145.71.34.2:255.255.255.0:SLIP uwalt:CSLIP,1006

     ttyS1::145.71.34.3:145.71.34.2:255.255.255.0:Dynamic ttyS1:CSLIP,296
      ..

The fields are:

1. login name: as returned by getpwuid(getuid()) or tty name.

2. unused: compat. with passwd

3. Remote Address: IP address of the calling host, either numeric or by name

4. Local Address: IP address of this machine, again numeric or by name

5. Netmask: in dotted decimal notation

6. Comment field: put whatever you want here.

7. protocol: Slip, CSlip etc.

8. MTU: decimal number

An example /etc/net/diphosts entry for a remote SLIP user might be:

     Sfredm::145.71.34.1:145.71.34.2:255.255.255.0:SLIP uwalt:SLIP,296

which specifies a SLIP link with remote address of 145.71.34.1 and MTU of 296, or:

     Sfredm::145.71.34.1:145.71.34.2:255.255.255.0:SLIP uwalt:CSLIP,1006

which specifies a cSLIP-capable link with remote address 145.71.34.1 and MTU of 1006.

Therefore, all users who you wish to be allowed a statically allocated dial-up IP access should have an entry in the /etc/diphosts. If you want users who call a particular port to have their details dynamically allocated then you must have an entry for the tty device and do not configure a user based entry. You should remember to configure at least one entry for each tty device that your dialup users use to ensure that a suitable configuration is available for them regardless of which modem they call in on.

When a user logs in they will receive a normal login and password prompt at which they should enter their SLIP-login userid and password. If these verify ok then the user will see no special messages and they should just change into SLIP mode at their end. The user should then be able to connect ok and be configured with the relevant parameters from the diphosts file.

6.28.3 SLIP server using the dSLIP package.

Matt Dillon <dillon@apollo.west.oic.com> has written a package that does not only dial-in but also dial-out SLIP. Matt’s package is a combination of small programs and scripts that manage your connections for you. You will need to have tcsh installed as at least one of the scripts requires it. Matt supplies a binary copy of the expect utility as it too is needed by one of the scripts. You will most likely need some experience with expect to get this package working to your liking, but don’t let that put you off.

Matt has written a good set of installation instructions in the README file, so I won’t bother repeating them.

You can get the dSLIP package from its home site at:

apollo.west.oic.com

     /pub/linux/dillon_src/dSLIP203.tgz

or from:

sunsite.unc.edu

     /pub/Linux/system/Network/serial/dSLIP203.tgz

Read the README file and create the /etc/passwd and /etc/group entries before doing a make install.

6.29 STRIP support (Starmode Radio IP)

STRIP device names are ‘st0’, ‘st1’, etc.

Kernel Compile Options:

     Network device support  --->
             [*] Network device support
             ....
             [*] Radio network interfaces
             < > STRIP (Metricom starmode radio IP)

STRIP is a protocol designed specifically for a range of Metricom radio modems for a research project being conducted by Stanford University called the MosquitoNet Project

<http://mosquitonet.Stanford.EDU/mosquitonet.html>. There is a lot of interesting reading here,

even if you aren’t directly interested in the project.

The Metricom radios connect to a serial port, employ spread spectrum technology and are typically capable of about 100kbps. Information on the Metricom radios is available from the: Metricom Web Server

<http://www.metricom.com/>.

At present the standard network tools and utilities do not support the STRIP driver, so you will have to download some customised tools from the MosquitoNet web server. Details on what software you need is

available at the: MosquitoNet STRIP Page <http://mosquitonet.Stanford.EDU/strip.html>.

A summary of configuration is that you use a modified slattach program to set the line discipline of a serial tty device to STRIP and then configure the resulting ‘st[0-9]’ device as you would for ethernet with one important exception, for technical reasons STRIP does not support the ARP protocol, so you must manually configure the ARP entries for each of the hosts on your subnet. This shouldn’t prove too onerous.

6.30 Token Ring

Token ring device names are ‘tr0’, ‘tr1’ etc. Token Ring is an IBM standard LAN protocol that avoids collisions by providing a mechanism that allows only one station on the LAN the right to transmit at a time. A ‘token’ is held by one station at a time and the station holding the token is the only station allowed to transmit. When it has transmitted its data it passes the token onto the next station. The token loops amongst all active stations, hence the name ‘Token Ring’.

Kernel Compile Options:

     Network device support  --->
             [*] Network device support
             ....
             [*] Token Ring driver support
             < > IBM Tropic chipset based adaptor support

Configuration of token ring is identical to that of ethernet with the exception of the network device name to configure.

6.31 X.25

X.25 is a circuit based packet switching protocol defined by the C.C.I.T.T. (a standards body recognised by Telecommunications companies in most parts of the world). An implementation of X.25 and LAPB are being worked on and recent 2.1.* kernels include the work in progress.

Jonathon Naylor jsn@cs.nott.ac.uk is leading the development and a mailing list has been established to discuss Linux X.25 related matters. To subscribe send a message to: majordomo@vger.rutgers.edu with the text ”subscribe linux-x25” in the body of the message.

Early versions of the configuration tools may be obtained from Jonathon’s ftp site at ftp.cs.nott.ac.uk

<ftp://ftp.cs.nott.ac.uk/jsn/>.

6.32 WaveLan Card

Wavelan device names are ‘eth0’, ‘eth1’, etc.

Kernel Compile Options:

     Network device support  --->
             [*] Network device support
             ....
             [*] Radio network interfaces
             ....
             <*> WaveLAN support

The WaveLAN card is a spread spectrum wireless lan card. The card looks very like an ethernet card in practice and is configured in much the same way.

You can get information on the Wavelan card from Wavelan.com <http://www.wavelan.com/>.

7 Cables and Cabling

Those of you handy with a soldering iron may want to build your own cables to interconnect two linux machines. The following cabling diagrams should assist you in this.

7.1 Serial NULL Modem cable

Not all NULL modem cables are alike. Many null modem cables do little more than trick your computer into thinking all the appropriate signals are present and swap transmit and receive data. This is ok but means that you must use software flow control (XON/XOFF) which is less efficient than hardware flow control. The following cable provides the best possible signalling between machines and allows you to use hardware (RTS/CTS) flow control.

     Pin Name  Pin                               Pin
     Tx Data    2  -----------------------------  3
     Rx Data    3  -----------------------------  2
     RTS        4  -----------------------------  5
     CTS        5  -----------------------------  4
     Ground     7  -----------------------------  7
     DTR        20 -\---------------------------  8
     DSR        6  -/
     RLSD/DCD   8  ---------------------------/-  20
                                              \-  6

7.2 Parallel port cable (PLIP cable)

If you intend to use the PLIP protocol between two machines then this cable will work for you irrespective of what sort of parallel ports you have installed.

     Pin Name    pin            pin
     STROBE      1*
     D0->ERROR   2  ----------- 15
     D1->SLCT    3  ----------- 13
     D2->PAPOUT  4  ----------- 12
     D3->ACK     5  ----------- 10
     D4->BUSY    6  ----------- 11
     D5          7*
     D6          8*
     D7          9*
     ACK->D3     10 ----------- 5
     BUSY->D4    11 ----------- 6
     PAPOUT->D2  12 ----------- 4
     SLCT->D1    13 ----------- 3
     FEED        14*
     ERROR->D0   15 ----------- 2
     INIT        16*

     SLCTIN      17*
     GROUND      25 ----------- 25

Notes:

• Do not connect the pins marked with an asterisk ‘*’.

• Extra grounds are 18,19,20,21,22,23 and 24.

• If the cable you are using has a metallic shield, it should be connected to the metallic DB-25 shell at one end only.

Warning: A miswired PLIP cable can destroy your controller card. Be very careful and double check every connection to ensure you don’t cause yourself any unnecessary work or heartache.

While you may be able to run PLIP cables for long distances, you should avoid it if you can. The specifications for the cable allow for a cable length of about 1 metre or so. Please be very careful when running long plip cables as sources of strong electromagnetic fields such as lightning, power lines and radio transmitters can interfere with and sometimes even damage your controller. If you really want to connect two of your computers over a large distance you really should be looking at obtaining a pair of thin-net ethernet cards and running some coaxial cable.

7.3 10base2 (thin coax) Ethernet Cabling

10base2 is an ethernet cabling standard that specifies the use of 52 ohm coaxial cable with a diameter of about 5 millimetres. There are a couple of important rules to remember when interconnecting machines with 10base2 cabling. The first is that you must use terminators at both ends of the cabling. A terminator is a 52 ohm resistor that helps to ensure that the signal is absorbed and not reflected when it reaches the end of the cable. Without a terminator at each end of the cabling you may find that the ethernet is unreliable or doesn’t work at all. Normally you’d use ‘T pieces’ to interconnect the machines, so that you end up with something that looks like:

      |==========T=============T=============T==========T==========|
                 |             |             |          |
                 |             |             |          |
               -----         -----         -----      -----
               |   |         |   |         |   |      |   |
               -----         -----         -----      -----

where the ‘|’ at either end represents a terminator, the ‘======’ represents a length of coaxial cable with BNC plugs at either end and the ‘T’ represents a ‘T piece’ connector. You should keep the length of cable between the ‘T piece’ and the actual ethernet card in the PC as short as possible, ideally the ‘T piece’ will be plugged directly into the ethernet card.

7.4 Twisted Pair Ethernet Cable

If you have only two twisted pair ethernet cards and you wish to connect them you do not require a hub. You can cable the two cards directly together. A diagram showing how to do this is included in the

Ethernet-HOWTO <Ethernet-HOWTO.html>

8 Glossary of Terms used in this document.

The following is a list of some of the most important terms used in this document.

ARP This is an acronym for the Address Resolution Protocol and this is how a network machine associates an IP Address with a hardware address.

ATM This is an acronym for Asynchronous Transfer Mode. An ATM network packages data into standard size blocks which it can convey efficiently from point to point. ATM is a circuit switched packet network technology.

client This is usually the piece of software at the end of a system where the user is. There are exceptions to this, for example, in the X11 window system it is actually the server with the user and the client runs on the remote machine. The client is the program or end of a system that is receiving the service provided by the server. In the case of peer to peer systems such as slip or ppp the client is taken to be the end that initiates the connection and the remote end, being called, is taken to be the server.

datagram A datagram is a discrete package of data and headers which contain addresses, which is the basic unit of transmission across an IP network. You might also hear this called a ‘packet’.

DLCI The DLCI is the Data Link Connection Identifier and is used to identify a unique virtual point to point connection via a Frame Relay network. The DLCI’s are normally assigned by the Frame Relay network provider.

Frame Relay Frame Relay is a network technology ideally suited to carrying traffic that is of bursty or sporadic nature. Network costs are reduced by having many Frame Relay customer sharing the same network capacity and relying on them wanting to make use of the network at slightly different times.

Hardware address This is a number that uniquely identifies a host in a physical network at the media access layer. Examples of this are Ethernet Addresses and AX.25 Addresses.

ISDN This is an acronym for Integrated Services Digital Network. ISDN provides a standardised means by which Telecommunications companies may deliver either voice or data information to a customers premises. Technically ISDN is a circuit switched data network.

ISP This is an acronym of Internet Service Provider. These are organisations or companies that provide people with network connectivity to the Internet.

IP address This is a number that uniquely identifies a TCP/IP host on the network. The address is 4 bytes long and is usually represented in what is called the ”dotted decimal notation”, where each byte is represented in decimal from with dots ‘.’ between them.

MSS The Maximum Segment Size (MSS) is the largest quantity of data that can be transmitted at one time. If you want to prevent local fragmentation MSS would equal MTU-IP header.

MTU The Maximum Transmission Unit (MTU ) is a parameter that determines the largest datagram than can be transmitted by an IP interface without it needing to be broken down into smaller units. The MTU should be larger than the largest datagram you wish to transmit unfragmented. Note, this only prevents fragmentation locally, some other link in the path may have a smaller MTU and the datagram will be fragmented there. Typical values are 1500 bytes for an ethernet interface, or 576 bytes for a SLIP interface.

route The route is the path that your datagrams take through the network to reach their destination.

server This is usually the piece of software or end of a system remote from the user. The server provides some service to one or many clients. Examples of servers include ftp, Networked File System, or Domain Name Server. In the case of peer to peer systems such as slip or ppp the server is taken to be the end of the link that is called and the end calling is taken to be the client.

window The window is the largest amount of data that the receiving end can accept at a given point in time.

9 Linux for an ISP ?

If you are interested in using Linux for ISP purposes the I recommend you take a look at the Linux ISP homepage <http://www.anime.net/linuxisp/> for a good list of pointers to information you might need and use.

10 Acknowledgements

I’d like to thank the following people for their contributions to this document (in no particular order): Axel Boldt, Arnt Gulbrandsen, Gary Allpike, Cees de Groot, Alan Cox, Jonathon Naylor, Claes Ensson, Ron Nessim, John Minack, Jean-Pierre Cocatrix, Erez Strauss.

A special thanks to Alessandro Rubini for his excellent feedback and contributed corrections.

11 Copyright.

The NET-3-HOWTO, information on how to install and configure networking support for Linux. Copyright (c) 1997 Terry Dawson.

This program is free software; you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.

This program is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU General Public License for more details.

You should have received a copy of the GNU General Public License along with this program; if not, write to the:

Free Software Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.